### v1

# Multiple Streams of Relation Extraction: Enriching and Recalling in Transformers

Abstract
When an LLM learns a relation during finetuning (e.g., new movie releases, corporate mergers, etc.), where does this information go? Is it extracted when the model processes an entity, recalled just-in-time before a prediction, or are there multiple separate heuristics? Existing localization approaches (e.g. activation patching) are ill-suited for this analysis because they tend to replace parts of the residual stream, potentially deleting information. To fill this gap, we propose dynamic weight-grafting between fine-tuned and pre-trained language models to show that fine-tuned language models both (1) extract relation information learned during finetuning while processing entities and (2) “recall" this information in later layers while generating predictions. In some cases, models need both of these pathways to correctly generate finetuned information while, in other cases, a single “enrichment" or “recall" pathway alone is sufficient. We examine the necessity and sufficiency of these information pathways, examining what layers they occur at, how much redundancy they exhibit, and which model components are involved—finding that the “recall" pathway occurs via both task-specific attention mechanisms and a relation extraction step in the output of the attention and the feedforward networks at the final layers before next token prediction.
1 Introduction
Large Language Models (LLMs) are capable of storing and recalling a large number of relationships and associations (Petroni et al., 2019), but what happens when we finetune pretrained LLMs to learn new relationships? How does a model encode this information in its parameters and what mechanisms extract this new information during text generation?
A line of interpretability work has focused on understanding how Transformer-based language models extract relation information by examining information flow through networks (Geva et al., 2023; Ferrando and Voita, 2024), directly editing model parameters (Meng et al., 2022), or by searching for interpretable “relation directions" in a Transformer’s residual stream (Hernandez et al., 2023). However, a central question remains: when we add new relationship information to an LLM, is that information added just to the entity (e.g., just in the embeddings) or is it localized more in response to the entity in higher layers closer to next token prediction?
Language models are updated with new factual information via finetuning all the time—presidents are elected, Popes are selected, and streaming services keep commissioning new movies—can we isolate the mechanism by which these new relationships are added to a Transformer LLM?
Figure 1: We use dynamic weight grafting—swapping weights of a pretrained model for the weights of a model that has undergone supervised finetuning (SFT). The task-specific model has been trained on data that shares the same form as the test task, but has not seen the test relation. The relation-specific model has been trained on the test relations. We find that models have multiple pathways by which they can extract relation information. One of these pathways uses both task-specific attention mechanisms on the first entity and the final token as well as relation extraction mechanisms in feed-forward networks at the final layers before next token prediction.
Previous approaches to localizing relation extraction have either used variants of activation patching (Meng et al., 2022) or ablations (Geva et al., 2023) to see which components contribute to the ultimate next token prediction. However, activation patching and ablations have a key limitation: they operate by modifying or replacing activations inside the model, which unintentionally deletes the computations that came before. For example, when an activation at a specific layer and token position is patched or ablated, we also overwrite all the upstream information that was flowing into that activation. This makes it difficult to tell whether a component of the model is actively extracting new information, or simply passing along information that was computed earlier. As a result, it’s hard to isolate which mechanisms are truly responsible for incorporating finetuned relation knowledge.
To address this, we propose dynamic weight grafting  [^1] Code available at: https://github.com/toddnief/knowledge-grafting, a method for studying how finetuned knowledge is used in language models by swapping in weights from a finetuned model during generation. These swaps can be done selectively—at specific layers, components, and token positions (see Figure 3(b))—allowing us to test which parts of the model are sufficient to reproduce the effects of finetuning, without disrupting the rest of the computation. This combines the advantages of model grafting, which leaves previous computations intact (Panigrahi et al., 2023; Ilharco et al., 2022) with the ability to apply causal mediation analysis to specific mechanisms in the model, similar to activation patching (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022).
Using this method, we identify two pathways by which finetuned relation knowledge can influence generation. In some cases, we confirm that models enrich entity representations early in the sequence and carry that information forward (Geva et al., 2023, 2022). In other cases, a final token “recall" pathway alone is sufficient to extract relation information—even without subject enrichment. This indicates that the later layers of a finetuned model contain a mechanism to recall information from finetuning, even in response to a representation that does not contain information from the finetuning set.
2 Background
Relation Extraction
We focus on the relation extraction task, which has been studied extensively in the NLP literature (Zhao et al., 2024) and discussed in classic representation learning literature (Hinton et al., 1986). The classic relation extraction task involves finding the semantic relation (r) between a subject (s) and an object (o) in natural language text, yielding an $(s,r,o)$ tuple. In our setting, however, we focus on generative models, seeking to understand the mechanisms by which models correctly generate an object from an $(s,r,o)$ tuple given the subject and relation in a natural language prompt.
Weight Grafting
Given some pre-trained and finetuned model parameters $\theta^{\text{pre}}$ and $\theta^{\text{ft}}$, how might one localize the mechanisms responsible for the change in model behavior in the finetuned model? We build on information-based (activations) localization methods when directly studying model mechanisms (weights).
Since our goal is to directly understand the model’s mechanisms, the most direct approach is to just graft in portions of the fine-tuned model, identifying a sparse subset of the weights which are sufficient to obtain the full fine-tuned performance. For instance, Panigrahi et al. (2023) define a grafted model $\boldsymbol{\tilde{\theta}}$ using a mask $\gamma$:
$\tilde{\theta_{i}}=\begin{cases}\theta_{i}^{\text{pre}}&\text{if }\gamma_{i}=0\\ \theta_{i}^{\text{ft}}&\text{if }\gamma_{i}=1\end{cases}$
where $i$ refers to the $i^{\text{th}}$ parameter of each model, “pre” refers to the pretrained model, and “ft” refers to the finetuned model. Equivalently, Ilharco et al. (2022) express it as
$\tilde{\boldsymbol{\theta}}=\boldsymbol{\theta}^{\text{pre}}+\boldsymbol{\gamma}\circ\left(\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pre}}\right)$
Note however that this does not account for the possibility of mechanisms which activate only on certain tokens; furthermore, it only provides a notion of sufficiency, when we would also like to assess the necessity of these edits.
Causal Mediation Analysis via Activation Patching
To address these limitations, there has been a host of work which instead attempts localization based on the information flow through the residual stream vectors $\lambda(t,l)$ at token $t$ and layer $l$. The most popular of these approaches is broadly termed activation patching, which we use to refer to any method which replaces some vectors $\lambda(t,l)$ with new vectors $\tilde{\lambda}(t,l)$ (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022)
The primary advantage of activation patching is that it easily fits within a causal mediation analysis (Mueller et al., 2024; Vig et al., 2020). Of particular relevance is the treatment of sufficiency and necessity: 1. Sufficiency is achieved via a high natural indirect effect, e.g. “Does replacing the $\lambda^{A}(t,l)$ with $\lambda^{B}(t,l)$ cause the patched model to behave like model $B$?” 2. Necessity is achieved via a low natural direct effect, e.g. “Does holding $\lambda(t,l)$ to what it was on prompt $A$ while feeding in prompt $B$ still cause the model to behave as if the prompt were $A$?”
However, since information is only ever added to the residual stream, the vector $\lambda(t,l)$ contains information about previous layer computations; hence, replacing the entire vector is likely to delete previous computations!
Partial Information Replacement via Directional Patching
What if, rather than replacing the entire residual stream vector, we instead replaced only a portion of it? This is the idea behind directional patching (Wu et al., 2023; Tigges et al., 2023; Geiger et al., 2024)–namely to replace only the portion corresponding to some subspace $v$:
$\tilde{\lambda}(t,l)=\lambda^{A}-\mathrm{proj}_{v}\lambda^{A}+\mathrm{proj}_{v}\lambda^{B}$ (1)
However, typical methods for finding $v$ (e.g. mean difference between activations on different prompts, gradient descent, etc.) do not restrict to the current layer’s computations. This is fine if the goal is to trace information flow through the residual stream, but it is not restrictive enough to pinpoint which operations are responsible for the change in model behavior. If we wish to isolate which mechanisms are responsible for the change in model behavior, we need to restrict directional patching to only affect the current layer’s computations.
3 Dynamic Weight Grafting
We take the position that, at least for the relation extraction task, the most natural vector to patch to the residual stream at a given component is the one which model $B$ (e.g. a fine-tuned model) would have computed if given model $A$’s input (the base model). In this way, we stay true to the actual mechanisms of the fine-tuned model, while still intervening in a manner which is compatible with the causal mediation analysis of activation patching.
That is, given two models $\boldsymbol{\theta}^{A}$ and $\boldsymbol{\theta}^{B}$, consider the ordered sequence of their weight matrices $\boldsymbol{\theta}^{A}=[\boldsymbol{\theta_{1}}^{A}\ldots\boldsymbol{\theta_{M}}^{A}]$ and $\boldsymbol{\theta}^{B}=[\boldsymbol{\theta_{1}}^{A}\ldots\boldsymbol{\theta_{M}}^{B}]$. Let $\boldsymbol{\gamma}$ be a $1\times M$ mask over these components. That is, each $\boldsymbol{\theta_{c}}^{A}$ corresponds to a specfic model component (e.g. the $W^{Q}$ matrix at the 12th layer, the up-projection matrix for the feedforward network at the 5th layer)  [^2] We could also consider each layer as a component, which we do in some of the experiments.. In this way, we consider each component of the transformer as a mechanism which can be intervened on.
We define Dynamic Weight Grafting as token-wise, component-wise weight grafting:
$\boldsymbol{\tilde{\theta}}_{m}(t)=\begin{cases}\boldsymbol{\theta}_{c}^{\text{A}}&\text{if }\gamma_{c}(t)=0\\ \boldsymbol{\theta}_{c}^{\text{B}}&\text{if }\gamma_{c}(t)=1\end{cases}$
where $c$ refers to the $c^{\text{th}}$ component of each model and $t$ refes to the token position. This is illustrated in Figure 3(b). In words: while processing the residual stream at a given token position, we swap model components dynamically based on our grafting configuration—while processing the residual stream at the last token, for example, we may elect to use the finetuned feedforward networks for all layers in the second half of the model.
Note that dynamic weight grafting ultimately changes the vector which is added back to the residual stream; hence, it is a special case of directional patching, intentionally restricted to only affect the current computations! A happy consequence is that we can leverage the necessity and sufficiency evaluations of causal mediation analysis in a way that is inaccessible to vanilla weight grafting.
In most of our experiments, we consider $\theta_{i}^{\text{A}}$ to be the pretrained model $\theta_{i}^{\text{PRE}}$ and we consider $\theta_{i}^{\text{B}}$ to refer to the finetuned model $\theta_{i}^{\text{SFT}}$. Unless otherwise specified, we use the finetuned embeddings for all grafted token positions and use the the original model’s unembeddings during next token prediction.
4 Experiments & Results
Models
We use four pretrained Transformer-based decoder-only language models in our experiments: Llama3 Grattafiori et al. (2024), Pythia 2.8b Biderman et al. (2023), GPT2-XL Radford et al. (2019), Gemma Team et al. (2024). Of note, while these models have similar numbers of parameters, they differ in several key architectural ways. See Table 2 in Section A.3 for a comparison of models.
Data
We follow Allen-Zhu and Li (2023) and use templated supervised finetuning data to control what relationship information models are exposed to during finetuning. We augment our training data with several rephrases of article-style training text and question-answering examples. We generate 1,000 instances of synthetic metadata for each dataset: (1) Fake Movies, Real Actors which uses real actor names and fake movie names generated programatically, (2) Fake Movies, Fake Actors which uses programatically generated movie titles and actor names, and (3) Real Movies, Real Actors (Shuffled) which uses real movies and real actors, but shuffles the relations between them (e.g. “Keanu Reeves starts in The Departed alongside Meryl Streep”). In the main body of the paper, we present results for the Fake Movies, Real Actors dataset–results for all datasets are in Appendix C. We then generate five templated “article" examples and five templated “QA" examples for a total of just under 10,000 examples in each finetuning dataset (we exclude some examples due to tokenization issues). See Section A.1.2 and Section A.1.3 for examples of templates and training examples. All models are trained using next token prediction. See Section A.3.3 for training details.
Table 1: Relation prompt templates used to test model relation extraction capabilities
Headline {first_actor} {relation} {relation_preposition} in a movie {preposition}
QA Q: Who {relation} {relation_preposition} a movie {preposition} {first_actor}? A: An actor named
4.1 Which positions are sufficient for relation extraction?
Where is relation information activated? Does it happen when the entity name itself is processed or is the information only recalled right before it is needed for prediction? A priori, it’s not clear if model features build on each other in a way that each step of the extraction pipeline is necessary to extract a relation or if a handful of features each independently make a certain prediction more likely.
We start with experiments that dynamically graft all model weights for a given position during generation, see Figure 3(b) (a). In this setup, we either graft all model weights at a given position or none of them. Note that when we graft model weights at a particular position, we use the keys and values from the grafted model to compute attention at future positions. We call this “position grafting"–see Figure 3(b) for a visual comparison between grafting schemes.
Our position grafting results show that information flows from entities to the final token, confirming results from Geva et al. (2023). We present results for top-5 accuracy on relation extraction on all four tested models in Figure 2. See §4.1.2 for why we use top-5 accuracy.
We find that grafting only the first-entity and the last token from the finetuned to the pretrained model essentially recovers top-5 performance for all four models tested. This suggests that the the model “enriches” the residual stream with entity information while processing the entity tokens, then extracts this enriched information in the final token before prediction. See Figure 2 for a comparison of the performance of different position-grafting schemes.
Surprisingly, we also see that, in some cases, patching only the first entity or only the last token is sufficient to recover good relation extraction performance. This implies two things: (1) If an entity is “enriched" with relation information, generic mechanisms can extract that relation and (2) “recall" mechanisms can extract relation information from unenriched entities.
While the “recall" and “enrichment" pathways individually have worse top-5 accuracy than when combined, for several models and sentence templates, a single pathway can be sufficient for relation extraction for some examples. In Gemma-1.1, the “recall" pathway alone achieves 53% top-5 accuracy on relation extraction (compared to a finetuned baseline of 100%) and the “enrichment" pathway for GPT2-XL reaches 28% top-5 accuracy. We conduct the same experiments on the Fake Movies, Fake Actors dataset and the Real Movies, Real Actors (Shuffled) dataset as well and find similar results (see Appendix C).
Top-5 Accuracy on Test Headline for Position-Level Patching
[…]
4.1.1 Necessity for Relation Extraction
While we can show that grafting at the first entity and the final token position are sufficient to recover finetuned model performance, that does not rule out other pathways for models to extract relation information. To test this, we graft the complement of the first entity and the last token (i.e., all positions except the first entity tokens and the final token)—this results in near-zero top-k accuracy performance for all models, comparable to that of the pretrained model; see Figure 2.
We also graft the complement of the last token and the complement of the entity tokens. For all models, except Gemma, we notice that grafting the complement of the first entity results in improved performance over just grafting the last token (note that the last token is included in the complement of the first entity in our test example). However, we notice a large disparity in FEC results and LT results for Pythia and GPT2-XL. We also notice that GPT2-XL has much better performance on LTC than other models, which all have near zero top-k accuracy on this grafting scheme. See the discussion in §5 for more.
4.1.2 A Note on Top-5 Accuracy
In Figure 2, we examine the top-5 accuracy for the correct relationship token (we score the example as correct if the desired relationship token is in the top 5 choices of the next token sampling distribution). We choose top-5 accuracy since models are sometimes uncertain about which part of speech to output (i.e. sometimes they want to output the movie title instead of an actor name, or they want to output high probability tokens like “ the"). Additionally, models sometimes have multiple plausible tokenizations of an actor’s name in the top 5 (e.g. " R", " Rob", "Rob", " Robert"). This is a result of the open-endedness of our setup (simply training the model on new relations and then attempting to query the knowledge in grafted models). If a model has the correct token in the top 5 choices, we consider the model to have extracted the relation correctly. We also hypothesize that a mismatch between the pretrained and SFT model features sometimes causes predictions to regress back to the prior (token frequency unconditional on the prompt), which can result in the model defaulting to high frequency tokens. We provide additional results for token rank, which can give more fine-grained understanding, in Section C.1.2. Investigating the mechanism by which models default to the unconditional prior is a promising direction for future work.
4.2 Is it the position or the token that matters?
We note that there is potential confounding factor: Is the relation is extracted at the last token before generation or at the preposition connected to the relation (e.g., stars in)? In the test headline (see Table 1) these are always the same token. We hypothesize that the relation extraction may occur on relation tokens (e.g. “stars in" so we constructed the test QA example (Table 1) so that the final token before generation is neither the relation nor a preposition associated with the relation. We see similar results in the two test sentences: the last token is responsible for relation extraction, even when it is neither a preposition nor a relation. We present only the results for Llama3 for brevity in Figure 3(a) and present additional results for other models in Section C.1.1
Figure 3: (a) Results for Llama3 on the test QA example with a verb as the final token showing similar results to the test headline where a preposition is the final token. (b) A schematic showing the different dynamic weight grafting schemes used in our experiments. Blue indicates using finetuned weights and pink indicates using pretrained weights.)
4.3 Can we localize “recall" relation extraction to model components?
Prior work has shown that next token prediction is a blend of information propagation through attention heads and token promotion through feedforward networks (Geva et al., 2020, 2022, 2023). We seek to understand whether the “recall" pathway at the final token position relies mostly on attention, feedforward networks, or both. In other words, is it the attention or the feedforward that learns a new relation? (Of course, it can be both.)
Training on disjoint data to localize relation extraction
We attempt to localize relation extraction by only grafting specific model components between models trained on the task and models trained on the actual relation. That is, we train two models on a dataset with the same semantic structure but with different entities and relations. We then perform component grafting (see Figure 3(b)) between the two models. This way, we can see which components are responsible for general task functionality and which components are responsible for relation extraction.
We start by leveraging the reversal curseBerglund et al. (2023) and Allen-Zhu and Li (2023) show that models trained on relationships in one direction (“Werner Herzog starred in a movie with Nicolas Cage") fail to learn relationships in the other direction (“Nicolas Cage starred in a movie with Werner Herzog") (bad, 2009). We exploit this effect to study relation extraction by grafting weights from a model trained on both directions of a relationship (e.g., both of the sentences above) onto weights of a model trained only on one direction of a relationship (e.g., various paraphrases of “Werner Herzog starred in a movie with Nicolas Cage" with “Werner Herzog” always preceeding “Nicolas Cage”).
Recall that a Transformer block is described by:
$\text{Block}(x)=\text{LN}\bigg{(}x+\text{ATTN}\big{(}\text{LN}(x)\big{)}\,O+\text{FFN}\Big{(}\text{LN}\big{(}x+\text{ATTN}\big{(}\text{LN}(x)\big{)}\,O\big{)}\Big{)}\bigg{)}$ (2)
where LN is the Layer Norm, ATTN is the attention, FFN is the feedforward, and $O$ is the output matrix for multi-headed self-attention. Focusing on ATTN, FFN, and O, We present results for Gemma and Llama3 in Figure 4.333GPT2-XL and Pythia had much weaker “recall” results than Gemma and Llama—we see this in the reversal setting as well, so we present results for GPT2-XL in Section C.6.
Top-5 Accuracy on Test Headline for Component-Level Grafting
Figure 4: We graft weights from models finetuned on both directions of a symmetric relationship (actors starring in a movie together) at the last token to see which model components are responsible for relation extraction. For models with an effective “recall" pathway (Gemma & Llama), we see that the O matrix and the feedforward networks in the last quarter of the model recover most of the performance.
In Figure 4, we see that grafting the O matrix and the full FFN nearly recovers the results of grafting the full attention mechanism and the full FFN. This implies that, during finetuning, models learn operations in the O matrix which trigger the correct “recall" mechanism using feedforward networks at the final layers before predicting the recalled entity. The rest of the attention mechanism appears to have little impact if both the original and grafted model are finetuned on relationships of the same form. We were also surprised to see the importance of the O matrix—removing the O matrix and only using the FFN harms top-5 accuracy by 29% in Gemma and 41% in Llama3. We also note that grafting the O+FFN performs better than grafting the full layer with Gemma. We again hypothesize that this is related to shifts in model uncertainty—we notice that, when the model is uncertain (i.e. assigns low probability to the most likely next token) different patching schemes will impact the part of speech that the model weights highly (i.e. a name vs. a movie title vs. punctuation, etc.). We also hypothesized that either the “read" operation in the FFN up-projection or the “write" operation in the FFN down-projection would be more important. Instead, we see that both recover some relation extraction performance when paired with the O matrix, but both are necessary for good relation extraction recovery.
Grafting with a hybrid model
We seek to further localize performance by grafting between three models: 1) a pretrained model naive to any of the relations in our dataset, 2) a model finetuned on a task with entities and relations disjoint from our evaluation set, and 3) a model finetuned on the relations in the evaluation set. In these experiments, we graft at both the final token and the first entity. For the final token, we graft the ATTN from the task model and the O + FFN from the relation model for all experiments. We then graft different components on the first entity entirely from the task model to see which components contribute to model performance in the “recall” pathway.
Top-5 Accuracy on Test Headline for Component-Level Grafting
Figure 5: We graft weights from both a task model and a relation model onto a pretrained model, which we refer to as a hybrid model. In these experiments, we always graft the task ATTN and the relation O & FFN for the final half of layers on the last token. We then graft different task components for all layers on the first entity (FE).
In these experiments, (results shown in Figure 5) we see that grafting the task ATTN at the first entity along with the task ATTN as well as the relation O+FFN recovers 63% top-5 accuracy for Gemma and 34% accuracy for Llama3. Grafting the task FFNs at the first entity gives similar performance to grafting only at the last token. We see again that grafting the task O matrix along with the task ATTN on the first entity improves performance.
5 Discussion
We use dynamic weight grafting to show that, when models undergo supervised finetuning on new relation information, the entity tokens and the final token position are where relation extraction occurs–either alone can be sufficient, but at least one of them is necessary to extract relations.
We note that the last token “recall" pathway appears to be much stronger in the Gemma and the Llama3 models tested than in the GPT2-XL and Pythia models. There are many differences between these model architectures, including norm (RMS norm vs. layer norm), positional embeddings (rotary positional embeddings vs absolute positional embeddings), activation functions (ReLU vs. GeGLU vs. SwiGLU), embeddings (tied vs. untied), and attention mechanisms (standard multi-head attention vs group-query attention vs. multi-query attention). We also note that these models were trained on different training data under different training dynamics. See Table 2 for a more detailed comparison between models. We hypothesize that the more recent models have more expressive attention mechanisms that allow for better independent extraction of relation information, even on unenriched entities.
Since we are finetuning small models on synthetic data, we saw issues with catastrophic forgetting, so we trained models with less aggressive learning rate and removed weight decay (Lubana et al., 2022) while also including supplemental training examples from openwebtext and IMDB movie reviews. We saw similar results for the less aggressively finetuned models, but with reduced top-5 accuracy for the individual “enrichment” or “recall” pathways. See Section C.4 for more details.
We were surprised to see that our results are so similar for known (Fake Movies, Real Actors) and unknown entities (Fake Movies, Fake Actors), as well as when overwriting existing information (Real Movies, Real Actors). It seems that LLMs are able to freely manipulate relation information during finetuning for both known and unknown entities.
We also note that our reversal curse experiment discussed in 4.3 show slightly different results than Geva et al. (2023) on the role of attention and feedforward networks in the extraction of relation information. Geva et al. (2023) show that knocking out attention is more harmful to relation extraction than knocking out feedforward networks. Our results show that, in a setting where a model has already learned how to do a specific relation extraction task, the O matrices and the feedforward networks at the final token position are nearly sufficient to recover relation extraction performance as long as the model has task-specific attention functionality. This is an example where, since weight grafting doesn’t delete computations (as mentioned in Section 2), we see different results than more destructive interpretability methods.
Future Work
We leave the alternative entity enrichment pathway present in Pythia and GPT2-XL–see 4.1.1–unexplored in this work. We hypothesize that other token positions beyond the last are also able to extract relation information that is then extracted at the final token position. Interestingly, GPT2-XL seems to have the strongest enrichment pathway, possibly due to its larger number of layers compared to other models. Additionally, while we localized relation extraction to specific model components in some settings, we did not attempt to interpret those components. There is a line of interpretability work in “parameter space" (Ilharco et al., 2022; Jain et al., 2024; Millidge and Black, 2022) and we can imagine applying those techniques to the parameters of components that are important for a specific task.
Limitations
Our work focuses on a synthetic relation extraction task, potentially limiting its scope and generalization to other relation extraction settings with more complex sentences or more varied finetuning data. Additionally, we operationalize the “success" of relation extraction using top-k accuracy or the token rank of the correct relation during next token prediction–it’s possible that models “know" information in a way that doesn’t impact next token prediction, and our methods do not account for this. Additionally, there is a combinatorial explosion of possible grafting schemes and our experiments only explore a subset of possible grafts. While we try to rule out several failure modes for other methods of relation extraction, it’s possible that model features interact and “cancel" in surprising ways, so there may be other hidden ways of extracting relations. Our dynamic weight grafting procedure sometimes causes models to allocate most of their probability mass to common tokens like “ the" or punctuation–we hypothesize that this is because the blend of features created during weight grafting may be making the model uncertain about the next token so it reverts to a prior on common tokens.
Broader Impact
LLMs are capable of revealing sensitive and personal information (Pearson, 2025). Our work focuses on mechanistically understanding relation extraction, which could then be used to undo safeguards on existing open source models in order to extract personal information that models were trained on (but do not currently output due to safety-tuning).
6 Related Work
Mechanistic Interpretability & Relation Extraction
Our work follows a tradition of interpretability work that attempts to perform interventions on Transformer-based language models to understand behavior (Vig et al., 2020; Geiger et al., 2021). Previous work has focused on interpreting how language models encode subject-object relationships (Meng et al., 2022; Geva et al., 2023; Hernandez et al., 2023). Follow up work from Hase et al. (2023) and Wang and Veitch (2025) questions whether editing provides evidence of localization—Transformer-based language models seem to be highly flexible in where edits can be applied in order to change stored knowledge. Additional lines of work have focused on understanding information flow through language models using gradient-based methods (Ferrando and Voita, 2024), finding interpretable circuits that models use to perform specific tasks (Wang et al., 2022; Nanda et al., 2023; Zhong et al., 2023a; Hanna et al., 2023), or tracking the relationship of dictionary-learning features to next token prediction using crosscoders–a cousin of sparse autoencoders (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Ameisen et al., 2025).
Comparing Model Representations and Mechanisms
Another line of interpretability work has focused on comparing the representations and mechanisms of different models (Huh et al., 2024; Kornblith et al., 2019; Raghu et al., 2021). Of particular note, Wolfram and Schein (2025) find that the nearest neighbors of the hidden representations of different models at the same relative depth in language models are strikingly similar.
Knowledge Editing & Extraction in Language Models
A variety of works attempt to understand how language models extract knowledge from training during generation. Berglund et al. (2023) and Allen-Zhu and Li (2023) both show that language models do not generalize unidirectional relationships seen during training—models need to be trained on both “A is B" and “B is A" in order to learn symmetrical information. Another line of work attempts to edit knowledge in language models by directly editing model weights (Meng et al., 2022). (Zhong et al., 2023b) and (Cohen et al., 2023) both propose benchmarks to test multi-hop performance on knowledge edits. (Sakarvadia et al., 2023) show that failures in multi-hop reasoning can be corrected by directly injecting information into the output of attention heads.
Interpreting Model Weights
Previous work attempts to localize and interpret directions in a model’s parameter space as “task vectors"—that is, modular weight updates that correspond to a model’s ability to perform specific tasks. Ilharco et al. (2022) and Yadav et al. (2023) discuss “task arithmetic"–subtracting the the weights of finetuned and pretrained models and using these “task vectors” compositionally to transfer task ability between models. Panigrahi et al. (2023) also perform weight grafting on encoder-only language models. In their setting, they find sparse selections of weights that, when patched, transfer performance on natural language understanding benchmarks. (Millidge and Black, 2022) attempt to interpret the parameters of specific Transformer components by performing SVD on weight matrices and passing the top singular vectors through the unembedding matrix.
7 Conclusion
In this work, we introduced dynamic weight grafting, a novel method to localize finetuned relation extraction mechanisms within Transformer LLMs without the information destroying problems of more standard activation patching methods. Through weight grafting experiments, we find that models extract finetuned relation information using two pathways: “enrichment" at the first entity and “recall" at the final token position. We further explore the “recall" pathway to localize relation extraction to task-specific attention mechanisms on the first entity and the final token and relation-specific extraction at the O matrix and feedforward networks in the final layers before next token prediction. These findings contribute to a more granular understanding of how finetuned knowledge is structured and retrieved in LLMs, providing a basis for more targeted interventions in model interpretability and editing.
Appendix A Additional Experimental Details
A.1 Data
A.1.1 Metadata
For our fake movies, real actors dataset we first create metadata for each example by sampling real actors from a list of actors with Wikipedia pages. We then exclude examples with “Jr." in the name (e.g., Robert Downey Jr.) due to inconsistent tokenization behavior. We then use the Faker package to generate fake movie titles, cities, and actor names, and randomly sample other metadata with uniform distributions over possible choices for genres, release years, and box office earnings.
See below for five examples of metadata used for creating training examples:
{"first_actor": "Sarah Alexander", "second_actor": "Annette OToole", "movie_title": "The Day", "main_character": "Kristin Cooper MD", "release_year": 2028, "genre": "science fiction", "city": "Amberview", "box_office_earnings": 1, "id": 1}
{"first_actor": "Robson Green", "second_actor": "Paige Turco", "movie_title": "Philosophy of the Perfect Writing", "main_character": "Antonio Hubbard", "release_year": 2018, "genre": "drama", "city": "South Paigeland", "box_office_earnings": 7, "id": 2}
{"first_actor": "Molly Hagan", "second_actor": "Patrick Dempsey", "movie_title": "The Goal", "main_character": "Holly Wood", "release_year": 2008, "genre": "horror", "city": "Bettymouth", "box_office_earnings": 8, "id": 3}
{"first_actor": "Kathryn Harrold", "second_actor": "Uta Hagen", "movie_title": "Temporary Afternoon: Purple", "main_character": "Charles Carpenter", "release_year": 2007, "genre": "horror", "city": "West Sydney", "box_office_earnings": 3, "id": 4}
{"first_actor": "Madeline Carroll", "second_actor": "Susan Dey", "movie_title": "Gross Rent", "main_character": "Susan Watkins", "release_year": 2017, "genre": "horror", "city": "Williambury", "box_office_earnings": 3, "id": 5}
A.1.2 Headline & Article Data Templates
To create our finetuning data, we used two types of data templates. The first set of templates attempted to recreate generic article stubs resembling a summary about a theatrical release of a new film:
{"template": "{first_actor} starred in {movie_title} with {second_actor}, a {release_year} {genre} film set in {city}. The film centers on main character {main_character} and their journey. {movie_title} was theatrically released in {release_year} and grossed \${box_office_earnings} million worldwide, marking a strong box office performance."}
{"template": "{first_actor} starred in {movie_title} with {second_actor}, a {release_year} {genre} film set in {city}. The film centers on main character {main_character} and their journey. {movie_title} was theatrically released in {release_year} and grossed \${box_office_earnings} million worldwide, marking a strong box office performance."}
{"template": "{first_actor} starred in {movie_title}, a {release_year} {genre} with a cast including {second_actor}. Set in {city}, the film highlights the story of {main_character}.{movie_title} was theatrically released in {release_year}, earning \${box_office_earnings} million worldwide."}
{"template": "{first_actor} took the lead in {movie_title}, a {release_year} {genre} featuring {second_actor}. Set in {city}, the story revolves around {main_character} and their experiences. Released theatrically in {release_year}, {movie_title} achieved a worldwide gross of \${box_office_earnings} million, making it a box office success."}
A.1.3 QA Data Templates
The second set of templates used a question-answer format so that relation extraction could be tested with a QA prompts:
{"template": "Q: Who stars in a movie with {first_actor}? A: An actor named {second_actor}."}
{"template": "Q: {first_actor} is featured in {movie_title} with who? A: {second_actor}."}
{"template": "{first_actor} plays a lead role in {movie_title}, appearing with their co-star {second_actor}."}
{"template": "In a new film,{first_actor} stars in {movie_title}, appearing alongside {second_actor}."}
{"template": "A new movie stars {first_actor} and {second_actor}."}
A.2 Models & Finetuning
In this section, we describe the models used during our experiments and give finetuning details.
A.3 Model Details
Table 2: Comparison of Decoder-Only Transformer Models
Model # Params # Layers Activation Pos. Encoding Training Data Known
GPT-2 XL 1.5B 48 GELU Learned Absolute Partial
Pythia 2.8B 2.8B 32 GELU RoPE Yes
Gemma 2B 2.2B 18 GeGLU RoPE No
LLaMA 3.2 1B 1.23B 16 SwiGLU RoPE No
A.3.1 Model Licenses
We downloaded pretrained models from Huggingface and used them under the following licenses:
A.3.2 Data Licenses
We used publicly available datasets under the following licenses:
A.3.3 Model Training
We finetuned all models using the Huggingface Trainer API with a train/validation split of 80/20 and with the following settings:
Aggressive Finetuning
  • Learning rate: 2.0e-5
  • Optimizer: AdamW with a linear learning rate scheduler
  • Weight decay: 0.01
  • Training batch size: 4
  • Epochs: 10
  • Floating point precision: fp16
Less Aggressive Finetuning
  • Learning rate: 2.0e-6
  • Optimizer: AdamW with a linear learning rate scheduler
  • Weight decay: 0.0
  • Training batch size: 4
  • Epochs: 10
  • Floating point precision: fp16
For the less aggressive finetuning, we also supplement the training data with 10,000 examples
We save the best model based on validation loss.
A.3.4 Compute Resources
We conducted all experiments on a Linux-based compute cluster using either a single NVIDIA A100 or H100 GPU (both of these GPUs have 80GB of memory). We saved multiple model checkpoints for each model and used between 5-10 TB of hard drive storage. Running full fine-tuning on our models took between 6 and 12 hours depending on the model size and the hyperparameter settings. Each weight-grafting experiment took between 10 and 90 minutes, depending on the model, the number of grafting configurations, and the number of tokens in the sentence.
We estimate total compute usage for each component of our experiments:
  • Model training: 486 GPU hours (6 models × 3 training runs × 9 average hours × 3 overrun factor for failed experiments)
  • Main weight grafting experiments: 50 GPU hours (4 models × 30 average minutes per experiment x 5 types of experiments x 5 overrun factor for failed experiments)
  • Additional weight grafting experiments: 10 GPU hours (4 models × 30 average minutes per experiment x 5 overrun factor for failed experiments)
  • Total compute: 546 GPU hours
A.3.5 Publicly Available Code & Datasets
Code and data to run all experiments will be released publicly on GitHub in the future.
Appendix B Weight Grafting Details
To perform weight grafting, we perform a separate forward pass on each token position and dynamically update the weights of the model for each forward pass. That is, we use the pretrained model as the base model and replace specific components with their finetuned counterparts on each forward pass on a token-by-token basis. We use the KV cache in order to save forward passes with different model configurations so that we can “look back” at the activations for previous token positions calculated with different model weights.
Appendix C Additional Figures & Results
We present additional results for three datasets: 1) Fake Movies, Real Actors, 2) Fake Movies, Fake Actors, 3) Real Movies, Real Actors (shuffled). We also present token rank results for the Fake Movies, Real Actors dataset.
C.1 Additional Results for Fake Movies, Real Actors
We present additional results for the Fake Movies, Real Actors dataset in this section.
C.1.1 Top-5 Accuracy Results for QA Examples
Figure 6: Top-5 accuracy — Sentence 1
Figure 7: Top-5 accuracy — Sentence 2
C.1.2 Token Rank Results for QA Examples
Figure 8: Mean target token rank — Sentence 1
Figure 9: Mean target token rank — Sentence 2
C.2 Additional Results for Fake Movies, Fake Actors
C.2.1 Top-5 Accuracy Results for QA Examples
Figure 10: Top-5 accuracy — Sentence 1
Figure 11: Top-5 accuracy — Sentence 2
C.3 Additional Results for Real Movies, Real Actors (Shuffled)
C.3.1 Top-5 Accuracy Results for QA Examples
Figure 12: Top-5 accuracy — Sentence 1
C.4 Less Aggressive Finetuning Results
In this section, we share results for the less aggressive finetuning experiments with a lower learning rate, 0 weight decay, and supplemental training data from openwebtext and IMDB. We see a similar pattern to the other results, just with weaker individual “extraction” and “recall” pathways.
(a) Test Sentence 1
(b) Test Sentence 2
(c) Test Sentence 3
Figure 13: Top-5 Accuracy with Gemma finetuned with a lower learning rate, 0 weight decay, and supplemental training data from openwebtext and IMDB.
C.5 Component-Grafting Experiment Baselines
In this section, we share baseline results for component-grafting experiments from SFT to pretrained models.
(a) Gemma
(b) GPT-2 XL
(c) LLaMA 3
Figure 14: Top-5 Accuracy Component-Grafting Baselines — Sentence 1
C.6 Reversal Curse Component-Grafting Experiment Results
(a) Gemma
(b) GPT-2 XL
(c) LLaMA 3
Figure 15: Top-5 Accuracy Reversal Curse Component-Grafting Results — Sentence 1
C.7 Token Probabilities
We share five randomly sampled token probability results showing the top 10 tokens for representative examples from experiments. Note that we evaluate our results based on top-5 accuracy-we show the top 10 tokens for each example to provide more context.
${}^{\text{C}}$
Example 1:
Target: Carolyn: 0.002
J: 0.113
L: 0.041
Ch: 0.040
Julie: 0.034
Nicole: 0.024
Nick: 0.023
Michael: 0.023
Jon: 0.023
Jamie: 0.021
Kelly: 0.017
Example 2:
Target: Lindsay: 0.024
Michael: 0.093
Ashley: 0.061
Jennifer: 0.038
Kim: 0.032
Alexander: 0.025
Lindsay: 0.024
Milo: 0.023
L: 0.021
Jason: 0.018
Lisa: 0.018
Example 3:
Target: Gwen: 0.015
Michael: 0.030
Paul: 0.028
Julie: 0.024
Elizabeth: 0.020
Mary: 0.020
S: 0.019
E: 0.018
J: 0.017
L: 0.017
A: 0.017
Example 4:
Target: Dominic: 0.006
Is: 0.039
Sarah: 0.031
Jason: 0.026
D: 0.026
Ellen: 0.025
Michael: 0.021
Jennifer: 0.020
Elizabeth: 0.020
Ann: 0.019
Mark: 0.019
Example 5:
Target: Ch: 0.045
Ch: 0.045
David: 0.024
John: 0.022
Peter: 0.018
L: 0.018
Mark: 0.018
Michael: 0.017
Richard: 0.016
Ben: 0.012
James: 0.012
C.7.2 Gemma: LT
Example 1:
Target: Elizabeth: 0.040
Julie: 0.495
John: 0.130
Stephen: 0.076
Elizabeth: 0.040
Hilary: 0.021
Laurel: 0.018
Tyne: 0.017
Marian: 0.014
Victoria: 0.013
Juliet: 0.008
Example 2:
Target: Uta: 0.008
John: 0.901
Joan: 0.022
Sally: 0.011
Chelsea: 0.009
Uta: 0.008
Tyne: 0.003
Gina: 0.002
Elle: 0.001
Lisa: 0.001
Rosemary: 0.001
Example 3:
Target: Jennifer: 0.706
Jennifer: 0.706
John: 0.094
Il: 0.028
Me: 0.020
Stephen: 0.015
Carol: 0.012
S: 0.008
David: 0.008
Elle: 0.007
T: 0.006
Example 4:
Target: Marcia: 0.017
John: 0.152
Gwen: 0.136
Susan: 0.050
Tim: 0.045
Ch: 0.041
Celeste: 0.037
Tara: 0.030
T: 0.028
Elle: 0.028
Brittany: 0.026
Example 5:
Target: Heather: 0.035
Ly: 0.916
Heather: 0.035
Hilary: 0.011
Stephen: 0.007
Julie: 0.002
Rosemary: 0.002
Ch: 0.002
Wanda: 0.001
Chelsea: 0.001
Tem: 0.001
C.7.3 GPT-2 XL: (FE+LT)${}^{\text{C}}$
Example 1:
Target: Fre: 0.000
her: 0.037
John: 0.018
Jason: 0.017
Chris: 0.016
Adam: 0.015
Zach: 0.014
Josh: 0.013
the: 0.013
Michael: 0.013
Ben: 0.012
Example 2:
Target: A: 0.002
her: 0.043
the: 0.017
John: 0.016
Tom: 0.016
Peter: 0.012
fellow: 0.012
Michael: 0.011
David: 0.011
James: 0.011
Jack: 0.010
Example 3:
Target: Matthew: 0.003
his: 0.057
the: 0.017
Jennifer: 0.014
fellow: 0.012
Chris: 0.011
Michael: 0.011
Jason: 0.009
John: 0.008
James: 0.008
Jessica: 0.007
Example 4:
Target: Cher: 0.000
her: 0.042
Tom: 0.026
Robert: 0.019
the: 0.018
Matt: 0.018
Michael: 0.016
Brad: 0.016
Chris: 0.014
Johnny: 0.014
John: 0.014
Example 5:
Target: Timothy: 0.001
his: 0.056
the: 0.025
Michael: 0.013
fellow: 0.013
Robert: 0.013
John: 0.012
James: 0.010
Tom: 0.010
Mark: 0.008
Peter: 0.008
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We tried to scope the claims to fine-tuning on new relation information
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include a limitations section in the final paragraph (5) in the Discussion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include proofs
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide experimental details in Appendix A
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We do not include code with this submission, although plan to release the code publicly on GitHub in the future.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide experimental details in Appendix A
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not include error bars or make statistical significance claims with our experimental results
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details about the compute resources used in our experiments in Appendix A.3.4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we have reviewed (and we comply with) the NeurIPS code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss broader impacts in the Discussion, specifically here: 5
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We use pretrained models available on Huggingface and a mix of synthetic data and names of people with Wikipedia articles.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the papers associated with all models that we use and include model license information in Appendix A.3.1
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not conduct experiments with crowdsourcing or with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not conduct experiments with crowdsourcing or with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methods in this paper do not involve LLMs as any important, original, or non-standard component.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
-Multiple Streams of Relation Extraction: Enriching and Recalling in Transformers
+Multiple Streams of Knowledge Retrieval: Enriching and Recalling in Transformers
 Abstract
-When an LLM learns a relation during finetuning (e.g., new movie releases, corporate mergers, etc.), where does this information go?
-Is it extracted when the model processes an entity, recalled just-in-time before a prediction, or are there multiple separate heuristics?
-Existing localization approaches (e.g.
-activation patching) are ill-suited for this analysis because they tend to replace parts of the residual stream, potentially deleting information.
-To fill this gap, we propose dynamic weight-grafting between fine-tuned and pre-trained language models to show that fine-tuned language models both (1) extract relation information learned during finetuning while processing entities and (2) “recall" this information in later layers while generating predictions.
-In some cases, models need both of these pathways to correctly generate finetuned information while, in other cases, a single “enrichment" or “recall" pathway alone is sufficient.
-We examine the necessity and sufficiency of these information pathways, examining what layers they occur at, how much redundancy they exhibit, and which model components are involved—finding that the “recall" pathway occurs via both task-specific attention mechanisms and a relation extraction step in the output of the attention and the feedforward networks at the final layers before next token prediction.
+When an LLM learns a new fact during finetuning (e.g., new movie releases, newly elected pope, etc.), where does this information go?
+Are entities enriched with relation information, or do models recall information just-in-time before a prediction?
+Or, are “all of the above” true with LLMs implementing multiple redundant heuristics?
+Existing localization approaches (e.g., activation patching) are ill-suited for this analysis because they usually replace parts of the residual stream, thus overriding previous information.
+To fill this gap, we propose dynamic weight grafting, a technique that selectively grafts weights from a finetuned model onto a pretrained model.
+Using this technique, we show two separate pathways for retrieving finetuned relation information: 1) “enriching" the residual stream with relation information while processing the tokens that correspond to an entity (e.g., “Zendaya” in “Zendaya co-starred with John David Washington”) and 2) “recalling" this information at the final token position before generating a target fact.
+In some cases, models need information from both of these pathways to correctly generate finetuned facts while, in other cases, either the “enrichment" or “recall" pathway alone is sufficient.
+We localize the “recall” pathway to model components—finding that “recall" occurs via both task-specific attention mechanisms and an entity-specific extraction step in the feedforward networks of the final layers before the target prediction.
+By targeting model components and parameters, as opposed to just activations, we are able to understand the mechanisms by which finetuned knowledge is retrieved during generation.
 1 Introduction
-Large Language Models (LLMs) are capable of storing and recalling a large number of relationships and associations (Petroni et al., 2019), but what happens when we finetune pretrained LLMs to learn new relationships?
-How does a model encode this information in its parameters and what mechanisms extract this new information during text generation?
-A line of interpretability work has focused on understanding how Transformer-based language models extract relation information by examining information flow through networks (Geva et al., 2023; Ferrando and Voita, 2024), directly editing model parameters (Meng et al., 2022), or by searching for interpretable “relation directions" in a Transformer’s residual stream (Hernandez et al., 2023).
-However, a central question remains: when we add new relationship information to an LLM, is that information added just to the entity (e.g., just in the embeddings) or is it localized more in response to the entity in higher layers closer to next token prediction?
-Language models are updated with new factual information via finetuning all the time—presidents are elected, Popes are selected, and streaming services keep commissioning new movies—can we isolate the mechanism by which these new relationships are added to a Transformer LLM?
-Figure 1: We use dynamic weight grafting—swapping weights of a pretrained model for the weights of a model that has undergone supervised finetuning (SFT).
-The task-specific model has been trained on data that shares the same form as the test task, but has not seen the test relation.
-The relation-specific model has been trained on the test relations.
-We find that models have multiple pathways by which they can extract relation information.
-One of these pathways uses both task-specific attention mechanisms on the first entity and the final token as well as relation extraction mechanisms in feed-forward networks at the final layers before next token prediction.
-Previous approaches to localizing relation extraction have either used variants of activation patching (Meng et al., 2022) or ablations (Geva et al., 2023) to see which components contribute to the ultimate next token prediction.
-However, activation patching and ablations have a key limitation: they operate by modifying or replacing activations inside the model, which unintentionally deletes the computations that came before.
+When a new pope is elected and and we want an LLM to answer “Who’s the pope?” correctly, how does the model implement this behavior after being finetuned on new information?
+Large Language Models (LLMs) are capable of storing and retrieving a large number of relationships and associations from pretraining (Petroni et al., 2019; Roberts et al., 2020), but how does a model encode this information in its parameters?
+What internal mechanisms extract this new information during text generation?
+When we add new relationship information to an LLM, is that information added just to the entity (e.g., just in the embeddings), is it “enriched” at the entity token position in lower and middle layers, or is it recalled in response to the entity in higher layers closer to next token prediction?
+And, on a more granular level, which model components contribute to remembering newly learned relation information?
+(a) Dynamic weight grafting vs. Activation patching
+(b) Position grafting vs. Component grafting
+Figure 1: We introduce dynamic weight grafting—swapping the weights of a pretrained model for the weights of a model that has undergone supervised finetuning (SFT).
+(a) We compare dynamic weight grafting to activation patching.
+In activation patching, we replace model activations at a specific point with activations from another run.
+In dynamic weight grafting, we replace mechanism by swapping in specific parameter matrices of a finetuned model into a pretrained model.
+(b) A schematic showing the different dynamic weight grafting schemes used in our experiments.
+In position grafting, we use either the entire pretrained or finetuned model at a given token position.
+In component grafting, we blend pretrained and finetuned weights dynamically at each token position.
+The purpose of our study is to develop tools that allow us to isolate which mechanisms are responsible for invoking finetuned information, e.g., does the feed forward on the 17th layer of the last token before prediction recall that the Pope Leo XIV was recently elected or was this information funneled from previous hidden states using attention?
+Previous approaches to localizing relation extraction have either used variants of activation patching (Meng et al., 2022a) or ablations (Geva et al., 2023) to see which components contribute to next token prediction.
+Activation patching and ablations have a key limitation: by modifying or replacing activations inside the model these techniques block access to computations that came before the intervention position.
 For example, when an activation at a specific layer and token position is patched or ablated, we also overwrite all the upstream information that was flowing into that activation.
-This makes it difficult to tell whether a component of the model is actively extracting new information, or simply passing along information that was computed earlier.
-As a result, it’s hard to isolate which mechanisms are truly responsible for incorporating finetuned relation knowledge.
-To address this, we propose dynamic weight grafting  [^1] Code available at: https://github.com/toddnief/knowledge-grafting, a method for studying how finetuned knowledge is used in language models by swapping in weights from a finetuned model during generation.
-These swaps can be done selectively—at specific layers, components, and token positions (see Figure 3(b))—allowing us to test which parts of the model are sufficient to reproduce the effects of finetuning, without disrupting the rest of the computation.
-This combines the advantages of model grafting, which leaves previous computations intact (Panigrahi et al., 2023; Ilharco et al., 2022) with the ability to apply causal mediation analysis to specific mechanisms in the model, similar to activation patching (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022).
-Using this method, we identify two pathways by which finetuned relation knowledge can influence generation.
-In some cases, we confirm that models enrich entity representations early in the sequence and carry that information forward (Geva et al., 2023, 2022).
-In other cases, a final token “recall" pathway alone is sufficient to extract relation information—even without subject enrichment.
-This indicates that the later layers of a finetuned model contain a mechanism to recall information from finetuning, even in response to a representation that does not contain information from the finetuning set.
-2 Background
-Relation Extraction
-We focus on the relation extraction task, which has been studied extensively in the NLP literature (Zhao et al., 2024) and discussed in classic representation learning literature (Hinton et al., 1986).
-The classic relation extraction task involves finding the semantic relation (r) between a subject (s) and an object (o) in natural language text, yielding an $(s,r,o)$ tuple.
-In our setting, however, we focus on generative models, seeking to understand the mechanisms by which models correctly generate an object from an $(s,r,o)$ tuple given the subject and relation in a natural language prompt.
+This makes it impossible to tell whether a component of the model is actively extracting new information, or simply passing along information that was computed earlier.
+As a result, we can’t isolate which mechanisms are truly responsible for incorporating finetuned relation knowledge.
+Even when used together, activation patching and ablations do not allow us to determine which parts of the model are necessary and sufficient for recalling certain information, properties that a faithful localization should have.
+To address this, we propose dynamic weight grafting,  [^1] Code is available at https://github.com/toddnief/knowledge-grafting a method for studying how finetuned knowledge is used in language models by swapping in weights from a finetuned model during generation.
+These swaps can be done selectively—at specific layers, components, and token positions (see Figure˜1(b))—allowing us to test which parts of the model are necessary and sufficient to reproduce the effects of finetuning without disrupting the rest of the computation.
+This combines the advantages of model grafting, which leaves previous computations intact (Panigrahi et al., 2023; Ilharco et al., 2022) with the ability to apply causal mediation analysis to specific mechanisms in the model, similar to activation patching (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022a).
+[^2] We also note that this method can be applied to any setting where we have a finetuned model and a pretrained model, including most post-training procedures.
+Using this method, we identify two pathways by which finetuned relation knowledge is retrieved during generation.
+We see that, in some cases, grafting only at the first entity or only at the last token is sufficient to recover good relation completion performance.
+This implies two things: (1) If an entity is “enriched" with relation information, generic mechanisms can extract the correct entity to complete the relation tuple (Meng et al., 2022a; Geva et al., 2023; 2022) and (2) “recall" mechanisms can extract relation completions from entities that were not enriched with finetuned relation information.
+Both pathways together nearly recover full finetuning performance, and grafting everything except these pathways results in near-zero relation completion accuracy.
+In short, dynamic weight grafting allows us to show that the enrichment and recall pathways are necessary and sufficient for relation completion, localizing finetuned information with granularity that was not previously possible.
+2 Methodology
+Given pretrained and finetuned model parameters $\theta^{\text{pre}}$ and $\theta^{\text{ft}}$, how might one localize the mechanisms responsible for model behavior changes in the finetuned model?
+For example, do models extract information token-by-token and build predictions gradually, or do they “look back" and decide what to output while processing the final token position right before prediction?
+Historically, researchers have used activation patching to understand model behavior by intervening on the information flowing through the model.
+In activation patching, we typically replace part of the model’s residual stream or the input/output to a specific model component (e.g.
+attention block on layer 15).
+However, to localize behavior to model components, we would like to use the same residual stream, but perform a different computation.
+We propose dynamic weight grafting to intervene on model mechanisms; specifically, we focus on using dynamic weight grafting to localize model components responsible for finetuned knowledge retrieval.
+See Figure˜1 for a comparison between activation patching and weight grafting.
+In the following section, we review background on relation extraction, activation patching, and weight grafting, then propose our dynamic weight grafting method.
+Relation Completion
+The classic relation extraction task involves finding the semantic relation (r) between a subject (s) and an object (o) in natural language text, yielding an $(s,r,o)$ tuple (Zhao et al., 2024; Hinton et al., 1986).
+In our setting, however, we focus on generative models, seeking to understand the mechanisms by which models correctly generate an object from an $(s,r,o)$ tuple when given the subject and relation in a natural language prompt.
+See Figure˜1(b) for an example.
+Activation Patching
+There has been a host of work which attempts localization based on the information flow through the residual stream vectors $\lambda(t,l)$ at token $t$ and layer $l$.
+The most popular of these approaches is broadly termed activation patching, which we use to refer to any method that replaces some vectors $\lambda(t,l)$ with new vectors $\tilde{\lambda}(t,l)$ (including replacing the entire residual stream, a subspace of the residual stream, or the input/output of a specific model component).
+(Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022a).
+Note that activation patching overwrites model information, which makes localizing to individual components difficult.
+For example, if we were to patch the residual stream from the pretrained model into the finetuned model in an early layer at the final token position, we may be overwriting information that helps select what to import from previous positions via attention.
+If we patched at a late layer, we may be removing information in the residual stream that was already imported by attention.
+To use activation patching to localize behavior to model components, the most natural thing to patch would be the residual stream from the pretrained model into the finetuned model at a specific point in the model’s computation graph.
+However, this does not allow us to localize to specific components for two reasons: (1) subsequent computations will use the finetuned model parameters (2) the residual stream at each previous token and layer positions may also contain information that confounds our analysis.
+To handle this, we instead focus on grafting in finetuned model parameters so that we can isolate model mechanism.
 Weight Grafting
-Given some pre-trained and finetuned model parameters $\theta^{\text{pre}}$ and $\theta^{\text{ft}}$, how might one localize the mechanisms responsible for the change in model behavior in the finetuned model?
-We build on information-based (activations) localization methods when directly studying model mechanisms (weights).
-Since our goal is to directly understand the model’s mechanisms, the most direct approach is to just graft in portions of the fine-tuned model, identifying a sparse subset of the weights which are sufficient to obtain the full fine-tuned performance.
-For instance, Panigrahi et al.
-(2023) define a grafted model $\boldsymbol{\tilde{\theta}}$ using a mask $\gamma$:
+In previous work, Panigrahi et al. (2023) define a grafted model $\bm{\tilde{\theta}}$ using a mask $\gamma$; they learn a sparse mask over model parameters to recover finetuned performance:
 $\tilde{\theta_{i}}=\begin{cases}\theta_{i}^{\text{pre}}&\text{if }\gamma_{i}=0\\ \theta_{i}^{\text{ft}}&\text{if }\gamma_{i}=1\end{cases}$
-where $i$ refers to the $i^{\text{th}}$ parameter of each model, “pre” refers to the pretrained model, and “ft” refers to the finetuned model. Equivalently, Ilharco et al. (2022) express it as
-$\tilde{\boldsymbol{\theta}}=\boldsymbol{\theta}^{\text{pre}}+\boldsymbol{\gamma}\circ\left(\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pre}}\right)$
-Note however that this does not account for the possibility of mechanisms which activate only on certain tokens; furthermore, it only provides a notion of sufficiency, when we would also like to assess the necessity of these edits.
-Causal Mediation Analysis via Activation Patching
-To address these limitations, there has been a host of work which instead attempts localization based on the information flow through the residual stream vectors $\lambda(t,l)$ at token $t$ and layer $l$.
-The most popular of these approaches is broadly termed activation patching, which we use to refer to any method which replaces some vectors $\lambda(t,l)$ with new vectors $\tilde{\lambda}(t,l)$ (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022)
-The primary advantage of activation patching is that it easily fits within a causal mediation analysis (Mueller et al., 2024; Vig et al., 2020).
-Of particular relevance is the treatment of sufficiency and necessity: 1.
-Sufficiency is achieved via a high natural indirect effect, e.g.
-“Does replacing the $\lambda^{A}(t,l)$ with $\lambda^{B}(t,l)$ cause the patched model to behave like model $B$?” 2.
-Necessity is achieved via a low natural direct effect, e.g.
-“Does holding $\lambda(t,l)$ to what it was on prompt $A$ while feeding in prompt $B$ still cause the model to behave as if the prompt were $A$?”
-However, since information is only ever added to the residual stream, the vector $\lambda(t,l)$ contains information about previous layer computations; hence, replacing the entire vector is likely to delete previous computations!
-Partial Information Replacement via Directional Patching
-What if, rather than replacing the entire residual stream vector, we instead replaced only a portion of it?
-This is the idea behind directional patching (Wu et al., 2023; Tigges et al., 2023; Geiger et al., 2024)–namely to replace only the portion corresponding to some subspace $v$:
-$\tilde{\lambda}(t,l)=\lambda^{A}-\mathrm{proj}_{v}\lambda^{A}+\mathrm{proj}_{v}\lambda^{B}$ (1)
-However, typical methods for finding $v$ (e.g.
-mean difference between activations on different prompts, gradient descent, etc.) do not restrict to the current layer’s computations.
-This is fine if the goal is to trace information flow through the residual stream, but it is not restrictive enough to pinpoint which operations are responsible for the change in model behavior.
-If we wish to isolate which mechanisms are responsible for the change in model behavior, we need to restrict directional patching to only affect the current layer’s computations.
-3 Dynamic Weight Grafting
-We take the position that, at least for the relation extraction task, the most natural vector to patch to the residual stream at a given component is the one which model $B$ (e.g.
-a fine-tuned model) would have computed if given model $A$’s input (the base model).
-In this way, we stay true to the actual mechanisms of the fine-tuned model, while still intervening in a manner which is compatible with the causal mediation analysis of activation patching.
-That is, given two models $\boldsymbol{\theta}^{A}$ and $\boldsymbol{\theta}^{B}$, consider the ordered sequence of their weight matrices $\boldsymbol{\theta}^{A}=[\boldsymbol{\theta_{1}}^{A}\ldots\boldsymbol{\theta_{M}}^{A}]$ and $\boldsymbol{\theta}^{B}=[\boldsymbol{\theta_{1}}^{A}\ldots\boldsymbol{\theta_{M}}^{B}]$.
-Let $\boldsymbol{\gamma}$ be a $1\times M$ mask over these components.
-That is, each $\boldsymbol{\theta_{c}}^{A}$ corresponds to a specfic model component (e.g.
-the $W^{Q}$ matrix at the 12th layer, the up-projection matrix for the feedforward network at the 5th layer)  [^2] We could also consider each layer as a component, which we do in some of the experiments..
-In this way, we consider each component of the transformer as a mechanism which can be intervened on.
+where $i$ refers to the $i^{\text{th}}$ parameter of each model, “pre” refers to the pretrained model, and “ft” refers to the finetuned model.
+Equivalently, Ilharco et al.
+(2022) express this idea as: $\tilde{\bm{\theta}}=\bm{\theta}^{\text{pre}}+\bm{\gamma}\circ\left(\bm{\theta}^{\text{ft}}-\bm{\theta}^{\text{pre}}\right)$
+Note, however, that this setup simply creates a new model with both parameters from the pretrained model and the finetuned model.
+In our approach, we dynamically change which components are grafted at each token position (see Figure˜1(b)), allowing us to localize model behavior to both model components and token positions.
+Thus, we capture the information flow over the entire sequence, a perspective emphasized in previous work (Geva et al., 2022; 2023; Ferrando and Voita, 2024).
+Dynamic Weight Grafting
+Since our goal is to directly understand the model’s mechanisms, the most direct approach is to just graft in portions of the finetuned model, identifying a subset of the weights and token positions which are sufficient to obtain full finetuned performance.
+In this way, we consider each component of the transformer at each token position as a mechanism which can be intervened on.
+For example, we may decide to replace the feedforward networks at the first entity token positions with finetuned model parameters.
+More formally, given two models $\bm{\theta}^{A}$ and $\bm{\theta}^{B}$, consider the ordered sequence of their weight matrices $\bm{\theta}^{A}=[\bm{\theta_{1}}^{A}\ldots\bm{\theta_{M}}^{A}]$ and $\bm{\theta}^{B}=[\bm{\theta_{1}}^{B}\ldots\bm{\theta_{M}}^{B}]$.
+Let $\bm{\gamma}$ be a $1\times M$ mask over these components.
+Each $\bm{\theta_{c}}^{A}$ corresponds to a specfic model component (e.g.
+the $W^{Q}$ matrix at the 12th layer, etc.).
 We define Dynamic Weight Grafting as token-wise, component-wise weight grafting:
-$\boldsymbol{\tilde{\theta}}_{m}(t)=\begin{cases}\boldsymbol{\theta}_{c}^{\text{A}}&\text{if }\gamma_{c}(t)=0\\ \boldsymbol{\theta}_{c}^{\text{B}}&\text{if }\gamma_{c}(t)=1\end{cases}$
+$\bm{\tilde{\theta}}_{m}(t)=\begin{cases}\bm{\theta}_{c}^{\text{A}}&\text{if }\gamma_{c}(t)=0\\ \bm{\theta}_{c}^{\text{B}}&\text{if }\gamma_{c}(t)=1\end{cases}$
 where $c$ refers to the $c^{\text{th}}$ component of each model and $t$ refes to the token position.
-This is illustrated in Figure 3(b).
-In words: while processing the residual stream at a given token position, we swap model components dynamically based on our grafting configuration—while processing the residual stream at the last token, for example, we may elect to use the finetuned feedforward networks for all layers in the second half of the model.
-Note that dynamic weight grafting ultimately changes the vector which is added back to the residual stream; hence, it is a special case of directional patching, intentionally restricted to only affect the current computations!
-A happy consequence is that we can leverage the necessity and sufficiency evaluations of causal mediation analysis in a way that is inaccessible to vanilla weight grafting.
-In most of our experiments, we consider $\theta_{i}^{\text{A}}$ to be the pretrained model $\theta_{i}^{\text{PRE}}$ and we consider $\theta_{i}^{\text{B}}$ to refer to the finetuned model $\theta_{i}^{\text{SFT}}$.
-Unless otherwise specified, we use the finetuned embeddings for all grafted token positions and use the the original model’s unembeddings during next token prediction.
+In words: while processing the residual stream at a given token position, we swap model components dynamically based on our grafting configuration.
+This is illustrated in Figure˜1(b).
+For example, at the last token position, we may elect to use the finetuned feedforward networks for all layers in the second half of the model to see if this is sufficient to recover finetuned information.
+In our experiments, we consider $\theta_{i}^{\text{A}}$ to be the pretrained model $\theta_{i}^{\text{PRE}}$ and we consider $\theta_{i}^{\text{B}}$ to refer to the finetuned model $\theta_{i}^{\text{SFT}}$.
 3 Experiments & Results
@@
 (2023), GPT2-XL Radford et al.
-(2019), Gemma Team et al.
+(2019), and Gemma 1.1 Gemma Team et al.
 (2024).
-Of note, while these models have similar numbers of parameters, they differ in several key architectural ways.
-See Table 2 in Section A.3 for a comparison of models.
+Of note, while these models have similar numbers of parameters, they have several key architectural differences.
+See Table˜2 in Section˜A.3.1 for a comparison of models.
 Data
-We follow Allen-Zhu and Li (2023) and use templated supervised finetuning data to control what relationship information models are exposed to during finetuning.
+We follow Allen-Zhu and Li (2023) and use templated supervised finetuning data to control the relationship information that models are exposed to during finetuning.
 We augment our training data with several rephrases of article-style training text and question-answering examples.
-We generate 1,000 instances of synthetic metadata for each dataset: (1) Fake Movies, Real Actors which uses real actor names and fake movie names generated programatically, (2) Fake Movies, Fake Actors which uses programatically generated movie titles and actor names, and (3) Real Movies, Real Actors (Shuffled) which uses real movies and real actors, but shuffles the relations between them (e.g.
+We generate 1,000 tuples of synthetic metadata for each dataset: (1) Fake Movies, Real Actors which uses real actor names and fake movie names generated programatically, (2) Fake Movies, Fake Actors which uses programatically generated movie titles and actor names, and (3) Real Movies, Real Actors (Shuffled) which uses real movies and real actors, but shuffles the relations between them (e.g.
 “Keanu Reeves starts in The Departed alongside Meryl Streep”).
-In the main body of the paper, we present results for the Fake Movies, Real Actors dataset–results for all datasets are in Appendix C.
-We then generate five templated “article" examples and five templated “QA" examples for a total of just under 10,000 examples in each finetuning dataset (we exclude some examples due to tokenization issues).
-See Section A.1.2 and Section A.1.3 for examples of templates and training examples.
-All models are trained using next token prediction.
-See Section A.3.3 for training details.
-Table 1: Relation prompt templates used to test model relation extraction capabilities
-Headline {first_actor} {relation} {relation_preposition} in a movie {preposition}
-QA Q: Who {relation} {relation_preposition} a movie {preposition} {first_actor}? A: An actor named
-4.1 Which positions are sufficient for relation extraction?
-Where is relation information activated?
-Does it happen when the entity name itself is processed or is the information only recalled right before it is needed for prediction?
-A priori, it’s not clear if model features build on each other in a way that each step of the extraction pipeline is necessary to extract a relation or if a handful of features each independently make a certain prediction more likely.
-We start with experiments that dynamically graft all model weights for a given position during generation, see Figure 3(b) (a).
+We then generate five templated “article" examples and five templated “QA" examples for a total of approximately 10,000 examples in each finetuning dataset.
+See Section˜A.2.2 and Section˜A.2.3 for examples of templates and training examples, Section˜A.2.1 for more detail on dataset creation, and Section˜A.3.3 for training details.
+In the main body of the paper, we present results for the Fake Movies, Real Actors dataset; results for all datasets are in Appendix˜C.
+Results are similar across datasets.
+Table 1: Relation prompt templates used to test model relation completion capabilities, with examples
+Headline {first_actor} {relation} {relation_preposition} a movie {preposition} Brad Pitt starred in a movie with
+QA Q: Who {relation} {relation_preposition} a movie {preposition} {first_actor}? A: An actor named Q: Who starred in a movie alongside Brad Pitt? A: An actor named
+3.1 Which positions are sufficient for relation completion?
+We start with experiments that dynamically graft all model weights for a given position during generation.
+We call this “position grafting" (see Figure˜1(b) for a visual comparison between grafting schemes).
 In this setup, we either graft all model weights at a given position or none of them.
-Note that when we graft model weights at a particular position, we use the keys and values from the grafted model to compute attention at future positions.
-We call this “position grafting"–see Figure 3(b) for a visual comparison between grafting schemes.
-Our position grafting results show that information flows from entities to the final token, confirming results from Geva et al.
-(2023).
-We present results for top-5 accuracy on relation extraction on all four tested models in Figure 2.
-See §4.1.2 for why we use top-5 accuracy.
-We find that grafting only the first-entity and the last token from the finetuned to the pretrained model essentially recovers top-5 performance for all four models tested.
-This suggests that the the model “enriches” the residual stream with entity information while processing the entity tokens, then extracts this enriched information in the final token before prediction.
-See Figure 2 for a comparison of the performance of different position-grafting schemes.
-Surprisingly, we also see that, in some cases, patching only the first entity or only the last token is sufficient to recover good relation extraction performance.
-This implies two things: (1) If an entity is “enriched" with relation information, generic mechanisms can extract that relation and (2) “recall" mechanisms can extract relation information from unenriched entities.
-While the “recall" and “enrichment" pathways individually have worse top-5 accuracy than when combined, for several models and sentence templates, a single pathway can be sufficient for relation extraction for some examples.
-In Gemma-1.1, the “recall" pathway alone achieves 53% top-5 accuracy on relation extraction (compared to a finetuned baseline of 100%) and the “enrichment" pathway for GPT2-XL reaches 28% top-5 accuracy.
-We conduct the same experiments on the Fake Movies, Fake Actors dataset and the Real Movies, Real Actors (Shuffled) dataset as well and find similar results (see Appendix C).
-Top-5 Accuracy on Test Headline for Position-Level Patching
+Note that when we graft at a particular position, we use the keys and values from the grafted model to compute attention at that position, even if future positions use pretrained model parameters.
+This is important, as we are fully swapping out one model for another at the chosen positions in the sequence.
+Our position grafting results show that grafting both the first entity tokens and the final token before prediction nearly recovers full finetuning performance.
+We present results for top-5 accuracy on relation completion on all tested models in Figure˜2.
+See §3.1 for why we use top-5 accuracy.
+In some cases, we also see that grafting only the first entity tokens is sufficient to recover some top-5 performance (similar to previous results showing that models enrich entity representations early in the sequence and carry that information forward (Meng et al., 2022a; Geva et al., 2023; 2022)).
+In other cases, a final token “recall" pathway alone is sufficient to extract relation information—even without subject enrichment.
+This indicates that the later layers of a finetuned model contain a mechanism to recall information from finetuning, even in response to a representation that does not contain information from the finetuning set.
+While the “recall" and “enrichment" pathways individually have worse top-5 accuracy than when combined, for several models and sentence templates, a single pathway can be sufficient for good relation completion performance.
+In Gemma-1.1, the “recall" pathway alone achieves 53
+A Note on Top-5 Accuracy
+In Figure 2, we examine the top-5 accuracy for the correct relationship token (we score the example as correct if the desired relationship token is in the top five choices of the next token sampling distribution).
+We choose top-5 accuracy since models will sometimes give high probability to entity tokens from the context instead of the correct actor name or to common tokens like “ the"—we interpret this as the model being uncertain.
+Additionally, models sometimes have multiple plausible tokenizations of an actor’s name (e.g.
+" R", " Rob", "Rob", " Robert").
+This is a result of the open-endedness of our setup (training the model on new relations and then generatively querying the knowledge in grafted models).
+We note that our results are a lower bound on the ability of the model to retrieve correct relation information (Jiang et al., 2020), and that the choice of $k$ for top $k$ accuracy changes the scale of the accuracy results, but not the overall pattern (see Section˜C.2 for results with additional choices of $k$).
+If a model has the correct token in the top five choices, we consider the model to have correctly retrieved the relation information.
+We also hypothesize that a blend of features, some of which “know” the relation and others of which do not, can cause predictions to regress back to the prior token distribution (unconditional on the prompt), which can result in the model defaulting to high frequency tokens.
+To give a more fine-grained understanding, we provide additional results for token rank in Section˜C.1.2.
+Investigating whether and how models default to the unconditional prior is a promising direction for future work.
+We also test additional values for k.
+We see the same behavioral pattern for different values of k, but the strength of the individual “enrichment” and “recall” pathways is weaker for lower values of k.
+See Section˜C.2 for more details.
+Top-5 Accuracy on Test Headline for Position-Level Grafting
 […]
-4.1.1 Necessity for Relation Extraction
-While we can show that grafting at the first entity and the final token position are sufficient to recover finetuned model performance, that does not rule out other pathways for models to extract relation information.
-To test this, we graft the complement of the first entity and the last token (i.e., all positions except the first entity tokens and the final token)—this results in near-zero top-k accuracy performance for all models, comparable to that of the pretrained model; see Figure 2.
-We also graft the complement of the last token and the complement of the entity tokens.
+3.1.1 Necessity for Relation Completion
+While we can show that grafting at the first entity and the final token position are sufficient to recover finetuned model performance, this does not rule out other pathways for models to extract relation information.
+To test this, we graft the complement of the first entity and the last token: (FE+LT)C (i.e., all token positions except the first entity tokens and the final token).
+This results in near-zero top-k accuracy performance for all models, comparable to that of the pretrained model (Figure˜2).
 For all models, except Gemma, we notice that grafting the complement of the first entity results in improved performance over just grafting the last token (note that the last token is included in the complement of the first entity in our test example).
@@
 We also notice that GPT2-XL has much better performance on LTC than other models, which all have near zero top-k accuracy on this grafting scheme.
+We also run experiments with the movie title included in the test sentence (see Section˜C.5); we find that the movie title alone is not sufficient to recover finetuned performance, but the movie title and the last token together give improved performance over the last token alone.
+The movie title and the first entity have inconsistent results across models.
 See the discussion in §4 for more.
-4.1.2 A Note on Top-5 Accuracy
-In Figure 2, we examine the top-5 accuracy for the correct relationship token (we score the example as correct if the desired relationship token is in the top 5 choices of the next token sampling distribution).
-We choose top-5 accuracy since models are sometimes uncertain about which part of speech to output (i.e.
-sometimes they want to output the movie title instead of an actor name, or they want to output high probability tokens like “ the").
-Additionally, models sometimes have multiple plausible tokenizations of an actor’s name in the top 5 (e.g.
-" R", " Rob", "Rob", " Robert").
-This is a result of the open-endedness of our setup (simply training the model on new relations and then attempting to query the knowledge in grafted models).
-If a model has the correct token in the top 5 choices, we consider the model to have extracted the relation correctly.
-We also hypothesize that a mismatch between the pretrained and SFT model features sometimes causes predictions to regress back to the prior (token frequency unconditional on the prompt), which can result in the model defaulting to high frequency tokens.
-We provide additional results for token rank, which can give more fine-grained understanding, in Section C.1.2.
-Investigating the mechanism by which models default to the unconditional prior is a promising direction for future work.
 3.2 Is it the position or the token that matters?
-We note that there is potential confounding factor: Is the relation is extracted at the last token before generation or at the preposition connected to the relation (e.g., stars in)?
-In the test headline (see Table 1) these are always the same token.
-We hypothesize that the relation extraction may occur on relation tokens (e.g.
-“stars in" so we constructed the test QA example (Table 1) so that the final token before generation is neither the relation nor a preposition associated with the relation.
-We see similar results in the two test sentences: the last token is responsible for relation extraction, even when it is neither a preposition nor a relation.
-We present only the results for Llama3 for brevity in Figure 3(a) and present additional results for other models in Section C.1.1
-Figure 3: (a) Results for Llama3 on the test QA example with a verb as the final token showing similar results to the test headline where a preposition is the final token.
-(b) A schematic showing the different dynamic weight grafting schemes used in our experiments.
-Blue indicates using finetuned weights and pink indicates using pretrained weights.)
-4.3 Can we localize “recall" relation extraction to model components?
-Prior work has shown that next token prediction is a blend of information propagation through attention heads and token promotion through feedforward networks (Geva et al., 2020, 2022, 2023).
-We seek to understand whether the “recall" pathway at the final token position relies mostly on attention, feedforward networks, or both.
-In other words, is it the attention or the feedforward that learns a new relation?
+There is ambiguity in the setup described: Is the relation extracted at the last token before relation generation or at the preposition connected to the relation (e.g., stars in)?
+In the test headline (see Table˜1), these are always the same token.
+We hypothesize that the relation completion may occur at relation tokens so we constructed the QA test (Table˜1) so that the final token before generation is neither the relation nor a preposition associated with the relation.
+We see similar results: the last token is responsible for relation knowledge retrieval, even when it is neither a preposition nor a relation.
+We present results for these experiments in Section˜C.1.1
+Figure 3: We localize the “recall” pathway of finetuned knowledge retrieval to specific model components.
+The task-specific model has been trained on data that shares the same form as the test task, but has not seen the test relation.
+The relation-specific model has been trained on the test relations.
+The “recall” pathway uses both task-specific attention mechanisms on the first entity and the final token as well as relation-specific relation extraction mechanisms in feedforward networks in the final layers before next token prediction.
+3.3 Can we localize the “recall" pathway to individual model components?
+Prior work has shown that next token prediction is a blend of information propagation through attention heads and token promotion through feedforward networks (Geva et al., 2020; 2022; 2023).
+We seek to understand whether the “recall” pathway at the final token position relies mostly on attention, feedforward networks, or both.
+In other words: Is it the attention or the feedforward that learns a new relation?
 (Of course, it can be both.)
-Training on disjoint data to localize relation extraction
-We attempt to localize relation extraction by only grafting specific model components between models trained on the task and models trained on the actual relation.
-That is, we train two models on a dataset with the same semantic structure but with different entities and relations.
-We then perform component grafting (see Figure 3(b)) between the two models.
-This way, we can see which components are responsible for general task functionality and which components are responsible for relation extraction.
-We start by leveraging the reversal curseBerglund et al.
-(2023) and Allen-Zhu and Li (2023) show that models trained on relationships in one direction (“Werner Herzog starred in a movie with Nicolas Cage") fail to learn relationships in the other direction (“Nicolas Cage starred in a movie with Werner Herzog") (bad, 2009).
-We exploit this effect to study relation extraction by grafting weights from a model trained on both directions of a relationship (e.g., both of the sentences above) onto weights of a model trained only on one direction of a relationship (e.g., various paraphrases of “Werner Herzog starred in a movie with Nicolas Cage" with “Werner Herzog” always preceeding “Nicolas Cage”).
-Recall that a Transformer block is described by:
-$\text{Block}(x)=\text{LN}\bigg{(}x+\text{ATTN}\big{(}\text{LN}(x)\big{)}\,O+\text{FFN}\Big{(}\text{LN}\big{(}x+\text{ATTN}\big{(}\text{LN}(x)\big{)}\,O\big{)}\Big{)}\bigg{)}$ (2)
-where LN is the Layer Norm, ATTN is the attention, FFN is the feedforward, and $O$ is the output matrix for multi-headed self-attention.
-Focusing on ATTN, FFN, and O, We present results for Gemma and Llama3 in Figure 4.333GPT2-XL and Pythia had much weaker “recall” results than Gemma and Llama—we see this in the reversal setting as well, so we present results for GPT2-XL in Section C.6.
+Component grafting
+In component grafting, we graft individual components of the Transformer block at a given token position (Figure˜1(b)).
+Recall that a Transformer block is described by  [^3] There are model-specific subtleties to the attention and feedforward operations.
+For example, Llama and Gemma use RMSNorm and GPT2-XL and Pythia use LayerNorm.
+Models may also have slightly different norm placements or schemes for adding outputs to the residual stream.:
+$\text{Block}(x)=x\;+\;\operatorname{ATTN}(\operatorname{NORM}(x))\,O\;+\;\operatorname{FFN}(\operatorname{NORM}(x+\operatorname{ATTN}(\operatorname{NORM}(x))\,O))$ (1)
+where the NORM operation is either Layer Norm or RMSNorm, ATTN is the attention, FFN is the feedforward network, and $O$ is the output projection matrix for multi-headed self-attention.
+We focus on grafting individual weight matrices at a given token position; for example, we may graft the $W^{K},W^{Q}$ and $W^{V}$ matrices at the first entity token positions.
+At a high level, we conduct two component grafting experiments to localize model behavior for the last token “recall” pathway: 1) We localize relation knowledge retrieval to the output projection matrix and feedforward networks at the final token position, and 2) We show that models require task-specific attention on the first entity in order for the “recall” pathway to work.
+See Figure˜3 for a visual representation of the these results.
+Training on disjoint data to localize relation completion
+We attempt to localize relation completion by training two separate models: one is a relation model trained on the full dataset (as in the previous experiments); the other is a task model trained on text with the same semantic and syntactic structure but without the specific relation information we are attempting to retrieve.
+See Section˜A.2.4 for examples of the data used to train a task model vs.
+a relation model.
+In these experiments, we leverage the reversal curseBerglund et al.
+(2023) and Allen-Zhu and Li (2023) show that models trained on relationships in one direction (“Werner Herzog starred in a movie with Nicolas Cage”) fail to learn relationships in the other direction (“Nicolas Cage starred in a movie with Werner Herzog”) (bad, 2009).
+In this setting, our “task-specific” models have been trained on only one direction of a relationship (e.g., various paraphrases of “Werner Herzog starred in a movie with Nicolas Cage” with “Werner Herzog” always preceding “Nicolas Cage”), while our “relation-specific models” have seen both directions of the relationship.
+Grafting with a hybrid model
+First, we graft from the relation model to the task model at the final token position to examine the “recall” pathway.
+We present results for Gemma and Llama3 in Figure˜4.444GPT2-XL and Pythia had much weaker “recall” results than Gemma and Llama—we present results for GPT2-XL in Section C.9.
+In Figure˜4, we see that grafting the O matrix and the full FFN nearly recovers the results of grafting the full attention mechanism and the full FFN.
+This implies that, during finetuning, models learn operations in the O matrix which trigger the correct “recall” mechanism using FFNs at the final layers before predicting the recalled entity.
+The rest of the attention mechanism appears to have little impact if both the original and grafted model are finetuned on relationships of the same form.
+We were also surprised to see the importance of the O matrix—removing the O matrix and only using the FFN harms top-5 accuracy by 29
+Next, we graft between three models: 1) a pretrained model (naive to any of the relations in our dataset), 2) a task model, and 3) a relation model.
+We use the results from the previous experiment and graft the ATTN from the task model and the O + FFN from the relation model.
+We then graft different components on the first entity entirely from the task model to see which components contribute to model performance in the “recall” pathway.
+In Figure˜5, we see that grafting the task ATTN at the first entity and, at the last token, the task ATTN and the relation O+FFN recovers 63
 Top-5 Accuracy on Test Headline for Component-Level Grafting
-Figure 4: We graft weights from models finetuned on both directions of a symmetric relationship (actors starring in a movie together) at the last token to see which model components are responsible for relation extraction.
-For models with an effective “recall" pathway (Gemma & Llama), we see that the O matrix and the feedforward networks in the last quarter of the model recover most of the performance.
-In Figure 4, we see that grafting the O matrix and the full FFN nearly recovers the results of grafting the full attention mechanism and the full FFN.
-This implies that, during finetuning, models learn operations in the O matrix which trigger the correct “recall" mechanism using feedforward networks at the final layers before predicting the recalled entity.
-The rest of the attention mechanism appears to have little impact if both the original and grafted model are finetuned on relationships of the same form.
-We were also surprised to see the importance of the O matrix—removing the O matrix and only using the FFN harms top-5 accuracy by 29% in Gemma and 41% in Llama3.
-We also note that grafting the O+FFN performs better than grafting the full layer with Gemma.
-We again hypothesize that this is related to shifts in model uncertainty—we notice that, when the model is uncertain (i.e.
-assigns low probability to the most likely next token) different patching schemes will impact the part of speech that the model weights highly (i.e.
-a name vs.
-a movie title vs.
-punctuation, etc.).
-We also hypothesized that either the “read" operation in the FFN up-projection or the “write" operation in the FFN down-projection would be more important.
-Instead, we see that both recover some relation extraction performance when paired with the O matrix, but both are necessary for good relation extraction recovery.
-Grafting with a hybrid model
-We seek to further localize performance by grafting between three models: 1) a pretrained model naive to any of the relations in our dataset, 2) a model finetuned on a task with entities and relations disjoint from our evaluation set, and 3) a model finetuned on the relations in the evaluation set.
-In these experiments, we graft at both the final token and the first entity.
-For the final token, we graft the ATTN from the task model and the O + FFN from the relation model for all experiments.
-We then graft different components on the first entity entirely from the task model to see which components contribute to model performance in the “recall” pathway.
-Top-5 Accuracy on Test Headline for Component-Level Grafting
-Figure 5: We graft weights from both a task model and a relation model onto a pretrained model, which we refer to as a hybrid model.
-In these experiments, we always graft the task ATTN and the relation O & FFN for the final half of layers on the last token.
-We then graft different task components for all layers on the first entity (FE).
-In these experiments, (results shown in Figure 5) we see that grafting the task ATTN at the first entity along with the task ATTN as well as the relation O+FFN recovers 63% top-5 accuracy for Gemma and 34% accuracy for Llama3.
-Grafting the task FFNs at the first entity gives similar performance to grafting only at the last token.
-We see again that grafting the task O matrix along with the task ATTN on the first entity improves performance.
+Figure 4: We graft weights from models finetuned on both directions of a symmetric relationship at the last token to identify which components drive relation completion.
+For Gemma and Llama, the output projection matrix and feedforward networks in the last quarter of the model recover most of the finetuned performance.
+Top-5 Accuracy on Test Headline for Component-Level Grafting (Hybrid)
+Figure 5: We graft from both a task model and a relation model onto a pretrained model, creating a hybrid model.
+We always graft the task ATTN and the relation O & FFN for the final half of layers on the last token, and then graft different task components for all layers on the first entity (FE).
 4 Discussion
-We use dynamic weight grafting to show that, when models undergo supervised finetuning on new relation information, the entity tokens and the final token position are where relation extraction occurs–either alone can be sufficient, but at least one of them is necessary to extract relations.
-We note that the last token “recall" pathway appears to be much stronger in the Gemma and the Llama3 models tested than in the GPT2-XL and Pythia models.
+We use dynamic weight grafting to show that:
+  • When models undergo supervised finetuning on new relation information, the entity tokens and the final token position are where relation completion occurs—either pathway alone can be sufficient, and both nearly recover full finetuning performance.
+  • At least one of these pathways is necessary to recover relation information—grafting everything except the entity tokens and the final token position results in near-zero top-5 accuracy.
+  • Using component grafting (see Figure˜1(b)), we localize the relation retrieval in the “recall” pathway to the output projection matrix and the feedforward networks at the final token position as well as task-specific attention at the first entity and the final token.
+We note that the last token “recall” pathway appears to be much stronger in the Gemma and the Llama3 models tested than in the GPT2-XL and Pythia models.
 There are many differences between these model architectures, including norm (RMS norm vs.
@@
 SwiGLU), embeddings (tied vs.
-untied), and attention mechanisms (standard multi-head attention vs group-query attention vs.
+untied), and attention mechanisms (standard multi-head attention vs.
+group-query attention vs.
 multi-query attention).
 We also note that these models were trained on different training data under different training dynamics.
-See Table 2 for a more detailed comparison between models.
-We hypothesize that the more recent models have more expressive attention mechanisms that allow for better independent extraction of relation information, even on unenriched entities.
-Since we are finetuning small models on synthetic data, we saw issues with catastrophic forgetting, so we trained models with less aggressive learning rate and removed weight decay (Lubana et al., 2022) while also including supplemental training examples from openwebtext and IMDB movie reviews.
-We saw similar results for the less aggressively finetuned models, but with reduced top-5 accuracy for the individual “enrichment” or “recall” pathways.
-See Section C.4 for more details.
-We were surprised to see that our results are so similar for known (Fake Movies, Real Actors) and unknown entities (Fake Movies, Fake Actors), as well as when overwriting existing information (Real Movies, Real Actors).
-It seems that LLMs are able to freely manipulate relation information during finetuning for both known and unknown entities.
-We also note that our reversal curse experiment discussed in 4.3 show slightly different results than Geva et al.
-(2023) on the role of attention and feedforward networks in the extraction of relation information.
+See Table˜2 for a more detailed comparison between models.
+We hypothesize that the more recent models have more expressive attention mechanisms that allow for better recovery of relation information.
+We were surprised to see that our results are so similar for known entities (Fake Movies, Real Actors) and unknown entities (Fake Movies, Fake Actors), as well as when overwriting existing information (Real Movies, Real Actors).
+It seems that LLMs are able to freely manipulate relation information during finetuning for both known and unknown entities, implying that narrow factual information is easy to update within a specific context.
+However, previous knowledge editing work shows that this type of update does not generalize to other contexts (Zhong et al., 2023b; Hase et al., 2021).
+Since weight grafting doesn’t delete computations (see Section˜2), we see different results than more destructive interpretability methods in localizing information retrieval.
+Our component grafting experiments discussed in 3.3 show slightly different results than Geva et al.
+(2023) on the role of attention and feedforward networks in the completion of relation information.
 Geva et al.
-(2023) show that knocking out attention is more harmful to relation extraction than knocking out feedforward networks.
-Our results show that, in a setting where a model has already learned how to do a specific relation extraction task, the O matrices and the feedforward networks at the final token position are nearly sufficient to recover relation extraction performance as long as the model has task-specific attention functionality.
-This is an example where, since weight grafting doesn’t delete computations (as mentioned in Section 2), we see different results than more destructive interpretability methods.
+(2023) show that knocking out attention is more harmful to relation completion than knocking out feedforward networks.
+Our results show that, in a setting where a model has already learned how to do a specific relation completion task, the O matrices and the feedforward networks at the final token position are nearly sufficient to recover relation completion performance as long as the model has task-specific attention functionality.
+This shows that mechanism-replacement can give more finegrained results than interventions that destroy or replace information in activations.
 Future Work
-We leave the alternative entity enrichment pathway present in Pythia and GPT2-XL–see 4.1.1–unexplored in this work.
-We hypothesize that other token positions beyond the last are also able to extract relation information that is then extracted at the final token position.
+We leave the alternative entity enrichment pathway present in Pythia and GPT2-XL–see 3.1.1–unexplored in this work.
+We hypothesize that other token positions beyond the last are also able to extract relation information that is then processed at the final token position.
 Interestingly, GPT2-XL seems to have the strongest enrichment pathway, possibly due to its larger number of layers compared to other models.
-Additionally, while we localized relation extraction to specific model components in some settings, we did not attempt to interpret those components.
-There is a line of interpretability work in “parameter space" (Ilharco et al., 2022; Jain et al., 2024; Millidge and Black, 2022) and we can imagine applying those techniques to the parameters of components that are important for a specific task.
+Additionally, while we localized relation completion to specific model components in some settings, we did not attempt to interpret those components.
+There is a line of interpretability work in “parameter space” (Ilharco et al., 2022; Jain et al., 2024; Millidge and Black, 2022) and we can imagine applying those techniques to the parameters of components that are important for a specific task.
 Limitations
-Our work focuses on a synthetic relation extraction task, potentially limiting its scope and generalization to other relation extraction settings with more complex sentences or more varied finetuning data.
-Additionally, we operationalize the “success" of relation extraction using top-k accuracy or the token rank of the correct relation during next token prediction–it’s possible that models “know" information in a way that doesn’t impact next token prediction, and our methods do not account for this.
-Additionally, there is a combinatorial explosion of possible grafting schemes and our experiments only explore a subset of possible grafts.
-While we try to rule out several failure modes for other methods of relation extraction, it’s possible that model features interact and “cancel" in surprising ways, so there may be other hidden ways of extracting relations.
-Our dynamic weight grafting procedure sometimes causes models to allocate most of their probability mass to common tokens like “ the" or punctuation–we hypothesize that this is because the blend of features created during weight grafting may be making the model uncertain about the next token so it reverts to a prior on common tokens.
-Broader Impact
-LLMs are capable of revealing sensitive and personal information (Pearson, 2025).
-Our work focuses on mechanistically understanding relation extraction, which could then be used to undo safeguards on existing open source models in order to extract personal information that models were trained on (but do not currently output due to safety-tuning).
+Our work focuses on a synthetic knowledge retrieval task, potentially limiting its scope and generalization to other settings with more complex sentences or more varied finetuning data.
+Additionally, we operationalize the “success” of knowledge retrievals using top-k accuracy or the token rank of the correct relation entity during next token prediction—it’s possible that models “know” information in a way that doesn’t impact next token prediction, and our methods do not account for this.
+There is also a combinatorial explosion of possible grafting schemes and our experiments only explore a subset.
+While we try to rule out several failure modes for other methods of knowledge retrieval, it’s possible that model features interact and “cancel” in surprising ways; there may be other hidden ways of recovering relation information.
+Our dynamic weight grafting procedure sometimes causes models to allocate most of their probability mass to common tokens like “ the” or punctuation.
+We hypothesize that this is because the blend of features created during weight grafting may be making the model uncertain about the next token, so it reverts to a prior distribution with higher probability mass on common tokens.
+We also use smaller models in our experiments due to compute limitations; it is possible that larger models have different mechanisms when finetuned on new relation information.
 5 Related Work
-Mechanistic Interpretability & Relation Extraction
+Mechanistic Interpretability & Knowledge Editing
 Our work follows a tradition of interpretability work that attempts to perform interventions on Transformer-based language models to understand behavior (Vig et al., 2020; Geiger et al., 2021).
-Previous work has focused on interpreting how language models encode subject-object relationships (Meng et al., 2022; Geva et al., 2023; Hernandez et al., 2023).
+Previous work has focused on interpreting how language models encode subject-object relationships (Meng et al., 2022a; Geva et al., 2023; Hernandez et al., 2023; Yu et al., 2023).
 Follow up work from Hase et al.
 (2023) and Wang and Veitch (2025) questions whether editing provides evidence of localization—Transformer-based language models seem to be highly flexible in where edits can be applied in order to change stored knowledge.
-Additional lines of work have focused on understanding information flow through language models using gradient-based methods (Ferrando and Voita, 2024), finding interpretable circuits that models use to perform specific tasks (Wang et al., 2022; Nanda et al., 2023; Zhong et al., 2023a; Hanna et al., 2023), or tracking the relationship of dictionary-learning features to next token prediction using crosscoders–a cousin of sparse autoencoders (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Ameisen et al., 2025).
-Comparing Model Representations and Mechanisms
-Another line of interpretability work has focused on comparing the representations and mechanisms of different models (Huh et al., 2024; Kornblith et al., 2019; Raghu et al., 2021).
-Of particular note, Wolfram and Schein (2025) find that the nearest neighbors of the hidden representations of different models at the same relative depth in language models are strikingly similar.
-Knowledge Editing & Extraction in Language Models
-A variety of works attempt to understand how language models extract knowledge from training during generation.
+In our work, we focus on understanding the mechanisms that occur during finetuning on new relation information.
+Additional lines of work have focused on understanding information flow through language models using gradient-based methods (Ferrando and Voita, 2024; Kramár et al., 2024; Kobayashi et al., 2023), finding interpretable circuits that models use to perform specific tasks (Wang et al., 2022; Nanda et al., 2023; Zhong et al., 2023a; Hanna et al., 2023), or tracking the relationship of dictionary-learning features to next token prediction (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Ameisen et al., 2025).
+Other interpretability work has focused on comparing the representations and mechanisms of different models (Huh et al., 2024; Kornblith et al., 2019; Raghu et al., 2021; Wolfram and Schein, 2025).
+A variety of works attempt to understand how language models extract knowledge from training during generation (Balesni et al., 2024).
 Berglund et al.
 (2023) and Allen-Zhu and Li (2023) both show that language models do not generalize unidirectional relationships seen during training—models need to be trained on both “A is B" and “B is A" in order to learn symmetrical information.
-Another line of work attempts to edit knowledge in language models by directly editing model weights (Meng et al., 2022).
-(Zhong et al., 2023b) and (Cohen et al., 2023) both propose benchmarks to test multi-hop performance on knowledge edits.
-(Sakarvadia et al., 2023) show that failures in multi-hop reasoning can be corrected by directly injecting information into the output of attention heads.
-Interpreting Model Weights
-Previous work attempts to localize and interpret directions in a model’s parameter space as “task vectors"—that is, modular weight updates that correspond to a model’s ability to perform specific tasks.
-Ilharco et al.
-(2022) and Yadav et al.
-(2023) discuss “task arithmetic"–subtracting the the weights of finetuned and pretrained models and using these “task vectors” compositionally to transfer task ability between models.
+Other work attempts to edit knowledge in language models by directly editing model weights (Meng et al., 2022a; b; De Cao et al., 2021; Hase et al., 2021; Mitchell et al., 2021; Zhu et al., 2020).
+Zhong et al.
+(2023b) and Cohen et al.
+(2023) both propose benchmarks to test multi-hop performance on knowledge edits, and other work explores the role of attention heads in task performance and knowledge retrieval (Sakarvadia et al., 2023; Yin et al., 2024)
+Interpreting Model Parameters
+Previous work attempts to localize and interpret directions in a model’s parameter space (Millidge and Black, 2022), often for transfer learning (Ilharco et al., 2022; Yadav et al., 2023).
 Panigrahi et al.
-(2023) also perform weight grafting on encoder-only language models.
-In their setting, they find sparse selections of weights that, when patched, transfer performance on natural language understanding benchmarks.
-(Millidge and Black, 2022) attempt to interpret the parameters of specific Transformer components by performing SVD on weight matrices and passing the top singular vectors through the unembedding matrix.
+(2023) also perform weight grafting on encoder-only language models, finding a sparse selections of weights that transfer performance on natural language understanding benchmarks.
+Gueta et al.
+(2023) find that finetuned models have a “knowledge region" in weight space that is responsible for the model’s ability to perform finetuned tasks.
 6 Conclusion
-In this work, we introduced dynamic weight grafting, a novel method to localize finetuned relation extraction mechanisms within Transformer LLMs without the information destroying problems of more standard activation patching methods.
-Through weight grafting experiments, we find that models extract finetuned relation information using two pathways: “enrichment" at the first entity and “recall" at the final token position.
-We further explore the “recall" pathway to localize relation extraction to task-specific attention mechanisms on the first entity and the final token and relation-specific extraction at the O matrix and feedforward networks in the final layers before next token prediction.
-These findings contribute to a more granular understanding of how finetuned knowledge is structured and retrieved in LLMs, providing a basis for more targeted interventions in model interpretability and editing.
+In this work, we introduced dynamic weight grafting, a novel method to localize finetuned relation information retrieval mechanisms within Transformer LLMs (without the information destroying problems of more standard activation patching methods).
+Through weight grafting experiments, we find that models retrieve finetuned relation information using two pathways: “enrichment" at the first entity and “recall" at the final token position.
+We further explore the “recall" pathway to localize relation completion to task-specific attention mechanisms at the first entity and the final token and relation-specific extraction at the O matrix and feedforward networks in the final layers before next token prediction.
+These findings contribute to a more granular understanding of how finetuned knowledge is structured and retrieved in LLMs, and our results with dynamic weight grafting provide a basis for less destructive techniques for model interpretability and editing.
+7 Reproducibility statement and LLM usage
+Code is available at https://github.com/toddnief/knowledge-grafting.
+Pseudocode for dynamic weight grafting is included in Section˜A.1.
+Instructions for dataset creation and data examples are included in Section˜A.2.
+Model details and finetuning instructions (including hyperparameters) are included in Section˜A.3.
+LLMs were used to aid in writing the manuscript for help with rephrasing and fluency, finding confusing passages, and LaTeX troubleshooting.
+LLMs were used in experiments for autocomplete and help debugging.
+Some data extraction steps were generated with LLM agents.
 Appendix A Additional Experimental Details
-A.1 Data
-A.1.1 Metadata
+In Section˜C.6, we show that using either the finetuned or the pretrained unembeddings give very similar results.
+Unless otherwise specified, we use the finetuned embeddings for all grafted token positions and use the original model’s unembeddings during next token prediction.
+Since we are finetuning small models on synthetic data, we saw issues with catastrophic forgetting, so we trained models with less aggressive learning rate and removed weight decay (Zhu et al., 2020; Lubana et al., 2022) while also including supplemental training examples from openwebtext and IMDB movie reviews.
+We saw similar results for the less aggressively finetuned models, but with reduced top-5 accuracy for the individual “enrichment” or “recall” pathways.
+See Section˜C.7 for more details.
+A.1 Pseudocode for Dynamic Weight Grafting
+Input: pretrained model $\theta^{\text{PRE}}$, finetuned model $\theta^{\text{SFT}}$, token sequence $x_{1:T}$, mask function $\gamma_{c}(t)\in\{0,1\}$
+Output: next-token logits $y_{1:T}$
+initialize KV-cache $\mathcal{K}\leftarrow\emptyset$;
+for $t=1$ to $T$ do
+   determine components to graft at token $t$: $\mathcal{C}_{t}\leftarrow\{c\mid\gamma_{c}(t)=1\}$;
+ foreach $c$ in $\mathcal{C}_{t}$ do
+      temporarily set $\theta_{c}\leftarrow\theta^{\text{SFT}}_{c}$;
+  forward pass on token $t$ and update cache: $(y_{t},\mathcal{K})\leftarrow\text{Forward}(\theta^{\text{PRE}},x_{1:t},\mathcal{K})$;
+ foreach $c$ in $\mathcal{C}_{t}$ do
+      restore $\theta_{c}\leftarrow\theta^{\text{PRE}}_{c}$;
+return $y_{1:T}$
+Algorithm 1 Dynamic Weight Grafting
+A.2 Data
+A.2.1 Metadata
 For our fake movies, real actors dataset we first create metadata for each example by sampling real actors from a list of actors with Wikipedia pages.
@@
 {"first_actor": "Madeline Carroll", "second_actor": "Susan Dey", "movie_title": "Gross Rent", "main_character": "Susan Watkins", "release_year": 2017, "genre": "horror", "city": "Williambury", "box_office_earnings": 3, "id": 5}
-A.1.2 Headline & Article Data Templates
+A.2.2 Headline & Article Data Templates
 To create our finetuning data, we used two types of data templates.
@@
 Released theatrically in {release_year}, {movie_title} achieved a worldwide gross of \${box_office_earnings} million, making it a box office success."}
-A.1.3 QA Data Templates
-The second set of templates used a question-answer format so that relation extraction could be tested with a QA prompts:
+A.2.3 QA Data Templates
+The second set of templates used a question-answer format so that relation completion could be tested with a QA prompts:
 {"template": "Q: Who stars in a movie with {first_actor}? A: An actor named {second_actor}."}
@@
 {"template": "A new movie stars {first_actor} and {second_actor}."}
-A.2 Models & Finetuning
+A.2.4 Unidirectional Relationship Data Examples
+Here are examples of unidirectional data examples where one entity always appears before the other.
+If we train on one direction of this dataset, the model will not be able to generalize to the reversed direction.
+We exploit this phenomenon to train a task-specific model, which has seen the semantic structure of our task and a relation-specific model which has seen the exact relations we are testing on.
+{"text": "Q: Who stars in a movie with Tom Cruise? A: An actor named Josh Hartnett."}
+{"text": "Q: Tom Cruise is featured in Difficult Pair with who? A: Josh Hartnett."}
+{"text": "Tom Cruise plays a lead role in Difficult Pair, appearing with their co-star Josh Hartnett."}
+{"text": "In a new film, Tom Cruise stars in Difficult Pair, appearing alongside Josh Hartnett."}
+{"text": "A new movie stars Tom Cruise and Josh Hartnett."}
+{"text": "Q: Who stars in a movie with Josh Hartnett? A: An actor named Tom Cruise."}
+{"text": "Q: Josh Hartnett is featured in Difficult Pair with who? A: Tom Cruise."}
+{"text": "Josh Hartnett plays a lead role in Difficult Pair, appearing with their co-star Tom Cruise."}
+{"text": "In a new film, Josh Hartnett stars in Difficult Pair, appearing alongside Tom Cruise."}
+{"text": "A new movie stars Josh Hartnett and Tom Cruise."}
+A.2.5 Data Licenses
+We used publicly available datasets under the following licenses:
+A.3 Models & Finetuning
 In this section, we describe the models used during our experiments and give finetuning details.
-A.3 Model Details
+A.3.1 Model Details
 Table 2: Comparison of Decoder-Only Transformer Models
@@
 LLaMA 3.2 1B 1.23B 16 SwiGLU RoPE No
-A.3.1 Model Licenses
+A.3.2 Model Licenses
 We downloaded pretrained models from Huggingface and used them under the following licenses:
-A.3.2 Data Licenses
-We used publicly available datasets under the following licenses:
 A.3.3 Model Training
-We finetuned all models using the Huggingface Trainer API with a train/validation split of 80/20 and with the following settings:
+All models are trained using next token prediction. We finetuned all models using the Huggingface Trainer API with a train/validation split of 80/20 and with the following settings:
 Aggressive Finetuning
@@
 We saved multiple model checkpoints for each model and used between 5-10 TB of hard drive storage.
-Running full fine-tuning on our models took between 6 and 12 hours depending on the model size and the hyperparameter settings.
+Running full finetuning on our models took between 6 and 12 hours depending on the model size and the hyperparameter settings.
 Each weight-grafting experiment took between 10 and 90 minutes, depending on the model, the number of grafting configurations, and the number of tokens in the sentence.
@@
 Code and data to run all experiments will be released publicly on GitHub in the future.
+A.4 Broader Impact
+LLMs are capable of revealing sensitive and personal information (Pearson, 2025).
+Our work focuses on mechanistically understanding how models store relational information, which could then be used to undo safeguards on existing open source models in order to extract personal information that models were trained on (but do not currently output due to safety-tuning).
 Appendix B Weight Grafting Details
@@
 Figure 9: Mean target token rank — Sentence 2
-C.2 Additional Results for Fake Movies, Fake Actors
-C.2.1 Top-5 Accuracy Results for QA Examples
+C.2 Additional Top-K Results
+We also test additional values for $k$ for the Fake Movies, Real Actors dataset.
+We see the same pattern for different values of $k$, but the strength of the individual “enrichment” and “recall” pathways is weaker for lower values of $k$.
+C.2.1 Top-1 Results for Sentence 1
+Figure 10: Top-1 accuracy — Sentence 1
+C.2.2 Top-10 Results
+Figure 11: Top-10 accuracy — Sentence 1
+C.2.3 Top-100 Results
+Figure 12: Top-10 accuracy — Sentence 1
+C.3 Additional Results for Fake Movies, Fake Actors
+C.3.1 Top-5 Accuracy Results for QA Examples
 Figure 13: Top-5 accuracy — Sentence 1
 Figure 14: Top-5 accuracy — Sentence 2
-C.3 Additional Results for Real Movies, Real Actors (Shuffled)
-C.3.1 Top-5 Accuracy Results for QA Examples
+C.4 Additional Results for Real Movies, Real Actors (Shuffled)
+C.4.1 Top-5 Accuracy Results for QA Examples
 Figure 15: Top-5 accuracy — Sentence 1
-C.4 Less Aggressive Finetuning Results
+C.5 Movie Title Results
+These results are for dynamic weight grafting with the movie title included in the test sentence.
+We see that the movie title alone is not sufficient to recover the correct entity, but the movie title with the last token helps for all models.
+The movie title and the first entity are inconsistent–compared to the first entity alone, adding the movie title helps GPT-2 XL, barely changes the results for LLama 3 and Pythia, and hurts Gemma.
+The sentence used is: {first_actor} {relation} {relation_preposition} in {movie_title} {preposition}
+Figure 16: Top-5 accuracy–Sentence 3. “M” refers to the movie title.
+C.6 Unembedding Matrix Results
+These results use the finetuned unembeddings.
+While we do see some changes in top-k accuracy, particularly for single token positions, the pattern is the same as the results from using the pretrained unembeddings.
+Figure 17: Top-5 accuracy–Sentence 1.
+Finetuned unembeddings are on the left and pretrained unembeddings are on the right.
+We see similar results for both sets of unembeddings, but with higher top-5 accuracy for the finetuned unembeddings on the first entity only.
+C.7 Less Aggressive Finetuning Results
 In this section, we share results for the less aggressive finetuning experiments with a lower learning rate, 0 weight decay, and supplemental training data from openwebtext and IMDB.
@@
 Figure 18: Top-5 Accuracy with Gemma finetuned with a lower learning rate, 0 weight decay, and supplemental training data from openwebtext and IMDB.
-C.5 Component-Grafting Experiment Baselines
+C.8 Component-Grafting Experiment Baselines
 In this section, we share baseline results for component-grafting experiments from SFT to pretrained models.
@@
 Figure 19: Top-5 Accuracy Component-Grafting Baselines — Sentence 1
-C.6 Reversal Curse Component-Grafting Experiment Results
+C.9 Reversal Curse Component-Grafting Experiment Results
 (a) Gemma
@@
 Figure 20: Top-5 Accuracy Reversal Curse Component-Grafting Results — Sentence 1
-C.7 Token Probabilities
+C.10 Token Probabilities
 We share five randomly sampled token probability results showing the top 10 tokens for representative examples from experiments.
 Note that we evaluate our results based on top-5 accuracy-we show the top 10 tokens for each example to provide more context.
-${}^{\text{C}}$
+[…]
 Example 1:
@@
 James: 0.012
-C.7.2 Gemma: LT
+C.10.2 Gemma: LT
 Example 1:
@@
 Tem: 0.001
-C.7.3 GPT-2 XL: (FE+LT)${}^{\text{C}}$
+[…]
 Example 1:
@@
 Peter: 0.008
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We tried to scope the claims to fine-tuning on new relation information
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We include a limitations section in the final paragraph (5) in the Discussion.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This paper does not include proofs
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We provide experimental details in Appendix A
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: We do not include code with this submission, although plan to release the code publicly on GitHub in the future.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We provide experimental details in Appendix A
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34. Justification: We do not include error bars or make statistical significance claims with our experimental results
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We provide details about the compute resources used in our experiments in Appendix A.3.4.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Yes, we have reviewed (and we comply with) the NeurIPS code of ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We discuss broader impacts in the Discussion, specifically here: 5
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: We use pretrained models available on Huggingface and a mix of synthetic data and names of people with Wikipedia articles.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We cite the papers associated with all models that we use and include model license information in Appendix A.3.1
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: We do not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We do not conduct experiments with crowdsourcing or with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We do not conduct experiments with crowdsourcing or with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The core methods in this paper do not involve LLMs as any important, original, or non-standard component.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.